### v1

# Convergent Privacy Framework with Contractive GNN Layers for Multi-hop Aggregations

Abstract.
Differential privacy (DP) has been integrated into graph neural networks (GNNs) to protect sensitive structural information, e.g., edges, nodes, and associated features across various applications. A common approach is to perturb the message-passing process, which forms the core of most GNN architectures. However, existing methods typically incur a privacy cost that grows linearly with the number of layers (Usenix Security’23), ultimately requiring excessive noise to maintain a reasonable privacy level. This limitation becomes particularly problematic when deep GNNs are necessary to capture complex and long-range interactions in graphs.
In this paper, we theoretically establish that the privacy budget can converge with respect to the number of layers by applying privacy amplification techniques to the message-passing process, exploiting the contractive properties inherent to standard GNN operations. Motivated by this analysis, we propose a simple yet effective Contractive Graph Layer (CGL) that ensures the contractiveness required for theoretical guarantees while preserving model utility. Our framework, $\mathsf{CARIBOU}$, supports both training and inference, equipped with a contractive aggregation module, a privacy allocation module, and a privacy auditing module. Experimental evaluations demonstrate that $\mathsf{CARIBOU}$ significantly improves the privacy-utility trade-off and achieves superior performance in privacy auditing tasks.
Differential Privacy, Convergent Privacy Analysis, Graph Neural Networks, Perturbed Message Passing, Privacy Amplification.
$\ast$ co-corresponding authors.
ccs: Security and privacy Privacy-preserving protocolsccs: Security and privacy Privacy protectionsccs: Computing methodologies Machine learning algorithms
1. Introduction
Graph neural networks (GNNs) (Kipf and Welling, [n. d.]; Gu et al., 2020; Velickovic et al., 2018), designed for operating over structural data, have achieved success in various domains, including social networks (Perozzi et al., 2014; Zhang et al., 2022b) and recommendation systems (Chen et al., 2024a; Wu et al., 2021b, 2023a). At their core, many GNN architectures are built upon the message-passing paradigm, where node representations are iteratively updated by aggregating information from their neighbors. However, graph structures often encode sensitive information about relationships and attributes. As a result, GNNs — like other machine learning models — are vulnerable to privacy attacks, including membership inference (Wu et al., 2022a; Ye et al., 2022; He et al., 2021b) and attribute inference (Zhang et al., 2022a; Sun et al., 2023). These vulnerabilities highlight the urgent need for robust privacy protection mechanisms in graph learning.
Differential privacy (DP) (Mironov, 2017; Wasserman and Zhou, 2010; Dong et al., 2022; Sun et al., 2024) has emerged as a foundational framework for providing formal guarantees against data leakage over the past two decades, with widespread applications in machine learning (Tang et al., 2023; Yu et al., 2019), synthetic data generation (Sun et al., 2024; Li et al., 2023a; McKenna et al., 2021; Patwa et al., 2023), and beyond. In the context of GNNs, recent works (Sajadmanesh and Gatica-Perez, 2020; Wu et al., 2023b) have advanced privacy protection through edge-level DP (EDP) (Wu et al., 2022b) and node-level DP (NDP) (Iftikhar and Wang, 2021; Day et al., 2016) guarantees. The primary approach employs perturbed message passing, which injects calibrated Gaussian or Laplace noise into aggregation layers to protect the edge or node memberships in the training graph. While these approaches have provided formal privacy guarantees, they share a critical limitation: the privacy loss grows linearly with the number of layers $K$ or graph hops. In other words, each additional aggregation layer compounds the privacy cost, ultimately requiring large amounts of noise to maintain a reasonable level of privacy protection. This, in turn, severely degrades model utility. This is particularly problematic when deep GNNs are needed for capturing complex relationships (Feng et al., 2022) or when analyzing graphs with long-range interactions (Dwivedi et al., 2022; Chen et al., 2024b; Wu et al., 2021a; Tönshoff et al., 2024; Thürlemann and Riniker, 2023), such as molecular graphs (Zhang et al., 2022c), as in these cases the message-passing mechanism requires multiple hops to reach distant nodes.
Interestingly, empirical studies (Wu et al., 2022a) have shown that membership inference attacks are not particularly more successful against deeper GNNs, suggesting that the linear dependency of privacy cost on network depth might be an overestimation. This observation aligns with the phenomenon known as “over-smoothing” (Chen et al., 2020a) in GNNs, where node representations become increasingly homogeneous as network depth increases and consequently making membership inference more challenging. This homogenization effect might actually provide inherent privacy benefits due to the contractive nature of GNN aggregation operations. This challenge motivates our central research question:
Can we achieve differentially private graph learning with a convergent (bounded) privacy budget, thereby improving the privacy-utility trade-off for deep GNNs?
Figure 1. Comparing Calibrated Noise for Perturbed Message Passing. Previous analysis requires $\sigma^{2}\propto O(K)$, and our analysis demands $\sigma^{2}\propto O\Bigl{(}(1-C_{\mathsf{L}}^{K})(1+C_{\mathsf{L}})/((1+C_{\mathsf{L}}^{K})(1-C_{\mathsf{L}}))\Bigl{)}$, where $C_{\mathsf{L}}$ is Lipchitz constant. With sensitivity constrained to norm $1$, the signal-to-noise ratio is markedly low as $K$ increases, severely impacting utility.
In this work, we answer this question affirmatively. Unlike previous studies, we propose a new privacy analysis that leverages the contractiveness of message-passing operations to achieve a convergent privacy budget with respect to the number of layers $K$ instead of a linear growth as in previous works; see Figure 1 for a comparison. We further design a novel privacy-preserving GNN framework, $\mathsf{CARIBOU}$, that incorporates this analysis into a practical model.
1.1. Overview of Convergent Privacy Analysis
For a perturbed message-passing GNN with $K$ layers, the standard approach analyzes the privacy loss at each step and then applies the DP composition theorem to aggregate the total privacy cost. This approach is common in existing privacy analyses of perturbed message-passing GNNs (Sajadmanesh and Gatica-Perez, 2020; Wu et al., 2023b; Chien et al., 2023), resulting in a privacy budget that scales linearly with $K$, specifically $\epsilon=O(K/\sigma^{2})+O(\sqrt{K/\sigma})$. Thus, as $K$ increases, the amount of injected noise $\sigma$ must also grow, leading to degraded utility — particularly when a small $\epsilon$ is desired; see Section 3 for a motivating empirical study.
Inspired by recent advances in privacy amplification (Chourasia et al., 2021; Bok et al., 2024) through hidden states and contractive iterative processes, we observe that a similar amplification effect can be exploited in GNNs. Here, contractiveness refers to the property that the distance between two inputs is reduced after applying the operation, implying reduced distinguishability of outputs from a privacy perspective. The potential privacy amplification in GNNs arises from the following two observations: (1) GNNs typically do not expose intermediate node embeddings during training or inference, focusing only on the final node representations. (2) Standard message-passing operations — such as those used in Graph Convolutional Networks (GCNs) (Kipf and Welling, [n. d.]) (the dominant model in practice and in empirical studies (Wu et al., 2022a)) — are inherently contractive, a property that also underlies the over-smoothing phenomenon (Chen et al., 2020a). We theoretically validate this insight by showing that the privacy loss of a $K$-layer perturbed message-passing process with contractive layers satisfies a convergent privacy budget. Specifically, instead of growing linearly with $K$, we show that the privacy budget follows a convergent form: $\epsilon=O\left(\frac{(1-C_{\mathsf{L}}^{K})(1+C_{\mathsf{L}})}{(1+C_{\mathsf{L}}^{K})(1-C_{\mathsf{L}})}\right),$ where $C_{\mathsf{L}}$ is the Lipschitz constant of the message-passing operator; see Theorem 6 for details. Our improved privacy analysis is achieved by recasting the multi-layer perturbed GNN as a Contractive Noisy Iteration (CNI) process (Bok et al., 2024) and applying the privacy convergence results established for CNIs.
To leverage this analysis in practice, we design a simple yet effective Contractive Graph Layer (CGL) that enforces the contractiveness required for our theoretical guarantees while maintaining model expressivity. The CGL layer builds upon standard GCN-style aggregation, augmented with residual connections (Chen et al., 2020b) and mean aggregation normalization (Alon and Yahav, 2020), ensuring expressiveness across many layers without exposing sensitive edge information.
We quantitatively characterize the privacy guarantees of CGL by carefully bounding the Lipchitz constant of the perturbed message-passing operation (Proposition 7) and the sensitivity of the perturbed message-passing with respect to both edge-level privacy (Theorem 8) and node-level privacy (Theorem 9). Together, these results allow us to explicitly quantify the privacy budget of the CGL layer using our general theory, culminating in the final privacy guarantee stated in Theorem 11.
1.2. $\mathsf{CARIBOU}$: Framework and Evaluation
Building on perturbed CGL, we realize a private framework $\mathsf{CARIBOU}$ for GNN inference and training. $\mathsf{CARIBOU}$ includes contractive aggregation module, privacy allocation module, and privacy auditing module. Together, these design choices enable us to achieve convergent privacy guarantees while maintaining strong GNN performance across graphs with varying interaction ranges.
To evaluate $\mathsf{CARIBOU}$, we conduct extensive experiments over nine graph datasets, including commonly-used real-world datasets and chain-structured datasets for developing configurable interaction ranges. The experimental results demonstrate that $\mathsf{CARIBOU}$ improves non-trivial utility over standard graph and chain-structured datasets. Compared with several SOTA baselines, $\mathsf{CARIBOU}$ shows significant utility improvements, especially in high privacy regimes, and reasonable computational overhead. Ablation studies are provided to understand the relation between privacy-utility hops and various ranges of graph, and the choice of hyper-parameters of CGL. In addition to privacy verification, we perform auditing experiments based on two membership inference attacks (Wu et al., 2022a; Olatunji et al., 2021b) , showing $\mathsf{CARIBOU}$’s resistance to attacks.
Contribution. To summarize, our contribution includes:
  1. (1)
    A novel privacy analysis for GNNs that leverages the contractiveness of message-passing operations to achieve convergent privacy costs, even for deep networks; (Section 5)
  2. (2)
    The design of perturbed CGL and a practical differentially private GNN framework – $\mathsf{CARIBOU}$ with provable privacy guarantees and superior utility-privacy tradeoffs; (Section 4)
  3. (3)
    Extensive experimental validation across multiple graph datasets with varying structural properties, demonstrating significant improvements over state-of-the-art private GNN approaches. (Section 6)
1.3. Related Works
DP for graphs.
Graph can be interpreted into the combination by edges and nodes. Corresponding to instance-level DP (Mironov, 2017; Abadi et al., 2016; Ye and Shokri, 2022), the “instance” of graph data can be an edge or a node, naturally called edge DP and node DP. Early stage works (Gehrke et al., 2011; Mir and Wright, 2009; Wang et al., 2013; Chen et al., 2014; Lu and Miklau, 2014) started by private statistics estimation or counting, or release of private graphs under edge DP. Node DP protects against revealing the presence or absence of an individual node along with all its adjacent edges. Prior works consider node-privacy (Kasiviswanathan et al., 2013; Iftikhar and Wang, 2021; Day et al., 2016; Xia et al., 2021; Imola et al., 2021) tailored to specific graph statistics constrained by bounded node degrees. Designing DP mechanisms for graphs requires thoughtful consideration of multi-hop aggregation, demanding careful design to balance privacy and utility.
Differentially private GNNs.
GNNs have emerged as a key approach for applications over graph-structured data, such as intrusion detection (Apruzzese et al., 2023; Sommer and Paxson, 2010), social recommendation (Xia et al., 2023), and drug discovery (Zhong and Mottin, 2023). Sharing finally established model can lead to privacy risks (Zhang et al., 2022a; Ye et al., 2022; He et al., 2021b), typically membership inference attack (MIA) (Carlini et al., 2022; Ye et al., 2022). MIA stems from “overfitting”, where models can memorize training memberships (Zhang et al., 2021). In GNNs, message-passing mechanisms can leak sensitive information about their edge- or node-level neighbors. To address these risks, existing research (Sajadmanesh and Gatica-Perez, 2020; Wu et al., 2023b; Chien et al., 2023; Daigavane et al., 2021; Olatunji et al., 2021a; Sajadmanesh et al., 2023) integrating DP with GNNs, advance the EDP or NDP by a perturbed message-passing mechanism that adds the calibrated Gaussian (or Laplace) noise to the message-passing layer. Albeit their success, leveraging the contractive hidden node embeddings in private GNNs for enhancing privacy analysis remains an underexplored avenue. More related works are detailed in Appendix F.
2. Preliminary
2.1. Message Passing Graph Neural Networks
Graph Neural Networks (GNNs) are a class of neural networks that operate on graph-structured data. Most GNNs follow the message-passing paradigm (Feng et al., 2022), where nodes iteratively aggregate information from their neighbors to update their representations.
Message Passing Layers.
Let ${\bm{G}}=({\bm{V}},{\bm{E}})$ be a graph, where ${\bm{V}}$ denotes the set of vertices (or nodes) and ${\bm{E}}$ denotes the set of edges. Let ${\bm{X}}^{(k)}\in\mathbb{R}^{|{\bm{V}}|\times d}$ be the node feature matrix at layer $k$, where $d$ is the dimension of the node features. Additionally, we use ${\bm{X}}_{u}^{(k)}\in\mathbb{R}^{d}$ to denote the feature vector of node $u$ at layer $k$. Each layer of a message passing GNN can be generally written as,
(1) $\text{MP}_{G}({\bm{X}}_{u}^{(k)}):=\sigma\left(\psi\left({\bm{X}}^{(k)},\oplus_{v\in{\mathcal{N}}(u)}\phi({\bm{X}}_{u}^{(k)},{\bm{X}}_{v}^{(k)})\right)\right),$
where $\sigma$ is a non-linear activation function, ${\mathcal{N}}(u)$ is the set of neighbors of node $u$, $\phi$ is a function that computes the message from node $v$ to node $u$, $\oplus$ represents the aggregation function that processes all messages from the neighbors of node $u$, and $\psi$ is a function that updates the node feature vector of node $u$ with the aggregated messages. Examples of message passing GNNs, such as GCN (Kipf and Welling, 2017) and its variant, are in Appendix A.
Applications of Message Passing GNNs.
Message passing GNNs leverage multiple layers to iteratively refine node representations, which are then employed in tasks like node classification (Kolluri et al., 2022), link prediction (Rao et al., 2024), and graph classification (Li et al., 2024). Long-range graphs (Dwivedi et al., 2022; Chen et al., 2024b; Tönshoff et al., 2024) enhance deeper GNNs (Thürlemann and Riniker, 2023; Wu et al., 2021a) by capturing dependencies between distant nodes, crucial for tasks like molecular property prediction (Ma et al., 2022), protein interaction modeling (Zhong et al., 2022), and model complex node interactions (Feng et al., 2022).
2.2. Differential Privacy for GNNs
Definition 0 (Differential Privacy (Dwork et al., 2006)).
Given a data universe $\mathcal{D}$, two datasets $D,D^{\prime}\subseteq\mathcal{D}$ are adjacent if they differ by only one data instance. A random mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private if for all adjacent datasets $D,D^{\prime}$ and for all events $S$ in the output space of $\mathcal{M}$, we have $\mathrm{Pr}(\mathcal{M}(D)\in S)\leq e^{\epsilon}\mathrm{Pr}(\mathcal{M}(D^{\prime})\in S)+\delta$.
Intuitively, DP (Dwork et al., 2006) theoretically quantifies the privacy of a model by measuring the indistinguishability of the outputs of a mechanism $\mathcal{M}$ on two adjacent datasets $D$ and $D^{\prime}$. It can be classified into bounded DP and unbounded DP depending on the construction of $D^{\prime}$, where the former is by replacing a data instance of $D$ and the latter is by addition / removal of a data sample of $D$. The privacy budget $\epsilon$ is smaller representing a stronger privacy guarantee, while $\delta$ is a slackness quantity that relaxes the pure DP constraint.
2.2.1. Privacy Definition on Graphs
In the context of graph data, the notion of adjacency refers to the graph structure, which can be defined as edge-level adjacency (Definition 2) and node-level adjacency (Definition 3).
Definition 0 (Edge-level adjacency (Karwa et al., 2011)).
Two graphs ${\bm{G}}_{1}=\{{\bm{V}}_{1},{\bm{E}}_{1}\}$ and ${\bm{G}}_{2}=\{{\bm{V}}_{2},{\bm{E}}_{2}\}$ are considered as edge-level neighboring if they differ in a single edge (through addition or removal of the edge), i.e., (${\bm{V}}_{2}={\bm{V}}_{1})\wedge(\lnot({\bm{E}}_{2}\cap{\bm{E}}_{1})=e_{i}$) where $e_{i}\in{\bm{E}}_{1}$.
Definition 0 (Node-level adjacency (Karwa et al., 2011)).
Two graphs ${\bm{G}}_{1}=\{{\bm{V}}_{1},{\bm{E}}_{1}\}$ and ${\bm{G}}_{2}=\{{\bm{V}}_{2},{\bm{E}}_{2}\}$ are considered as node-level neighboring if they differ in a single node and its incident edges (through addition or removal of the node and its incident edges), i.e., $\lnot({\bm{V}}_{2}\cap{\bm{V}}_{1})=\{n_{i},\{e_{ij}\}_{\forall j}\}$ where $n_{i}\in({\bm{V}}_{1}\cup{\bm{V}}_{2})$ and $\{e_{ij}\}_{\forall j}$ connects to $n_{i}$.
2.2.2. Perturbed Message Passing with DP
To incorporate differential privacy into GNNs, one can add noise to the message passing layer, following the perturbed message passing approach (Sajadmanesh et al., 2023). Given a graph ${\bm{G}}$ and message passing function $\mathrm{MP}$, we define a sequence $\{{\bm{X}}^{(k)}\}_{k=0}^{K}$ of node feature matrices by:
(2) ${\bm{X}}^{(k+1)}=\Pi_{\mathcal{K}}(\mathrm{MP}_{{\bm{G}}}({\bm{X}}^{(k)})+{\bm{Z}}^{(k)})$
where ${\bm{X}}^{(0)}={\bm{X}}$ is the input feature matrix, ${\bm{Z}}^{(k)}\sim\mathcal{N}(0,\sigma^{2})$ is Gaussian noise, and $\Pi_{\mathcal{K}}$ projects features back to bounded set $\mathcal{K}$ (typically constraining $\|{\bm{X}}_{u}\|_{2}\leq 1$ for each node $u$). The privacy guarantees of perturbed message passing depend on the sensitivity of the mechanism:
Definition 0 (Sensitivity of Perturbed Message Passing).
Let $\mathrm{MP}_{{\bm{G}}}$, $\mathrm{MP}_{{\bm{G}}^{\prime}}$ be the perturbed message passing mechanisms applied to neighboring graphs $G,G^{\prime}$. The sensitivity is defined as:
(3) $\Delta(\mathrm{MP})=\max_{G,G^{\prime}}\max_{{\bm{X}}\in\mathcal{K}}\|\mathrm{MP}_{{\bm{G}}}({\bm{X}})-\mathrm{MP}_{{\bm{G}}^{\prime}}({\bm{X}})\|_{F}$
where the maximum is taken over all adjacent graphs and all node feature matrices in $\mathcal{K}$.
The sensitivity determines the scale of noise required for privacy guarantees. Lower sensitivity allows for less noise addition while maintaining the same privacy level, directly affecting the utility-privacy trade-off in differentially private GNNs.
Privacy Accounting.
Privacy accounting process analyzes the total privacy budget for the composition of several (adaptive) private algorithms. A common approach for analyzing the Gaussian mechanism in perturbed message passing is through Rényi differential privacy (RDP) (Mironov, 2017) and its composition theorem.
Definition 0 (Rényi differential privacy (Mironov, 2017)).
A randomized algorithm $\mathcal{M}$ is $(\alpha,\epsilon)$-RDP for $\alpha>1$, $\epsilon>0$ if for every adjacent datasets $X,X^{\prime}$, we have $D_{\alpha}(\mathcal{M}(X)\|\mathcal{M}(X^{\prime}))\leq\epsilon$, where $D_{\alpha}(P\|Q)$ is the Rényi divergence of order $\alpha$ between probability distributions $P$ and $Q$ defined as:
$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\mathbb{E}_{x\sim Q}\left[\left(\frac{P(x)}{Q(x)}\right)^{\alpha}\right]$
Theorem 6 (Composition of RDP (Mironov, 2017)).
If $\mathcal{M}_{1},\dots,\mathcal{M}_{k}$ are randomized algorithms satisfying, respectively, $(\alpha,\epsilon_{1})$-RDP, …, $(\alpha,\epsilon_{k})$-RDP, then their composition defined as $(\mathcal{M}_{1}(S),\dots,\mathcal{M}_{k}(S))$ satisfies $(\alpha,\epsilon_{1}+\cdots+\epsilon_{k})$-RDP.
In this work, we present our privacy results in terms of RDP for ease of interpretation, while our underlying analysis employs the tighter $f$-DP framework. This analysis leverages recent advances in privacy amplification techniques (Schuchardt et al., 2024; Bok et al., 2024) to achieve stronger privacy guarantees. The technical details of our convergent privacy analysis are discussed in §3.2.
3. Multi-hop Private GNNs Initiative
Deeper GNNs (Thürlemann and Riniker, 2023; Wu et al., 2021a) are vital in tasks like modeling molecular structures, where some properties depend on long range interactions (Dwivedi et al., 2022; Chen et al., 2024b; Tönshoff et al., 2024) of the nodes (Zhang et al., 2022c; Li et al., 2023b). In this long-range scenario, message passing GNNs must flow across multiple hops before reaching a target node—requiring stacking multiple layers to exchange information across $K$-hop neighborhoods. However, ensuring privacy in multi-layer GNNs poses key challenges. In this section, we outline and illustrate these challenges via a motivating case study (Section 3.1), ultimately motivating our new design insights (Section 3.2) based on contractive message passing.
3.1. Observations on Privacy Accumulation and Performance Degradation
A Case Study on GAP’s Performance on Learning Long-Range Interactions.
To further investigate this phenomenon, we evaluate GAP on the chain-structured dataset (see Figure 2), which is adopted in Gu et al. (2020); Liu et al. (2021) to examine long-range interaction learning capabilities. Specifically, this dataset creates a controlled environment for evaluating GNN performance on long-range interactions, as an MLP model would necessarily fail since most nodes have zero-valued feature vectors regardless of their chain type. Consequently, for a message-passing GNN to correctly classify nodes positioned $K$ hops away from the informative first node, it must perform at least $K$ propagation steps to transfer the meaningful features across the chain as shown in Figure 2. This requirement becomes particularly challenging in the private setting as noise must be injected after each message passing layer, potentially overwhelming the signal being propagated.
Figure 2. Message Passing on Chain-structured Dataset
Figure 3 compares GAP’s non-private accuracy (blue solid line) with its private versions (dashed lines), under different privacy budgets $\epsilon\in\{1,2,4,8,16,32\}$. It shows a binary node classification over multiple $8$-node chains as a case study. The red solid line in Figure 3 represents random guessing (50% accuracy).
Figure 3. Motivating Experiments of Classification Model over Chain-structured Dataset.
Our exploration reveals a stark contrast between private and non-private settings:
  • Non-private setting. GAP’s accuracy consistently improves with increasing layer depth, as message passing enables feature propagation across the chain. The model achieves satisfactory performance after sufficient depth ($K\geq 8$), ultimately reaching perfect classification (100% accuracy) at $K=15$ hops—demonstrating the necessity of deep architectures for capturing long-range dependencies.
  • Private setting. Privacy protection dramatically degrades model utility. As shown in Figure 3, even with relatively generous privacy budgets ($\epsilon=16$ or $\epsilon=32$), performance remains marginally above random guessing (50%). Crucially, increasing depth offers no benefit and often harms performance, as noise accumulates exponentially across layers. This confirms our theoretical concerns: standard approaches to privacy in GNNs fundamentally limit the ability to learn long-range interactions.
Figure 4. Overview of $\mathsf{CARIBOU}$. The framework integrates three key modules: the Contractive Aggregation Module (CAM) for generating embeddings with adjustable coefficients, the Privacy Allocation Module (PAM) for efficient privacy budget allocation, and the Privacy Auditing Module (PDM) for empirical privacy verification. Combining CAM and PAM can form the Perturbed Contractive Message Passing (PCMP). Together, these modules enable scalable and accurate private GNN learning under a convergent privacy budget.
Empirical (Counter-Intuitive) Observation: Deeper GNNs May Enhance Privacy. Recent work by Wu et al. (2024) revealed a counter-intuitive phenomenon: deeper GNNs empirically exhibit lower vulnerability to membership and link inference attacks. This challenges standard privacy composition analysis, which suggests privacy risk increases with model depth due to multiple queries of the graph data. The insight stems from the over-smoothing [^1] Incidentally, another similar phrase is “over-fitting”, which refers to a model performing well on training data but poorly on unseen data due to memorization. We emphasize that they are different concepts to avoid conflation. (Chen et al., 2020a) phenomenon, where node representations become indistinguishably homogeneous phenomenon, where node representations converge toward similar values as depth increases, making it inherently difficult for adversaries to distinguish individual nodes or infer sensitive relationships.
This observation suggests that standard privacy analysis may be overly pessimistic. The conventional approach of adding noise that scales linearly with depth may be unnecessarily conservative, as the natural privacy amplification properties of over-smoothing could enable tighter privacy bounds with less noise per layer.
3.2. Our Approach: Contractive Message Passing for Convergent Privacy
Our key insight is to identify and leverage the inherent privacy amplification that occurs in deeper GNNs — a phenomenon that appears counter-intuitive but can be formally explained through the mathematical properties of message passing operations. This privacy amplification is intrinsically connected to the well-known over-smoothing phenomenon in GNN. Over-smoothing (Chen et al., 2020a) emerges fundamentally from the contractiveness of typical graph aggregation operations. When nodes aggregate information from their neighbors (e.g., through averaging as in GCN), the resulting representations necessarily become more similar to each other. This mathematical property can be formally defined:
Definition 0 (Contractive Map).
[…]
Designing Private GNNs with Convergent Privacy.
We aim to leverage the privacy amplification properties of contrative map to design a new framework for private GNNs. This framework is inspired by the recent advances in differentially private gradient descent (DP-GD) (Chourasia et al., 2021; Bok et al., 2024), which has shown that the privacy cost can converge to a finite value even with arbitrarily many iterations. Motivated by it, we aim to translate the advanced privacy analysis techniques from DP-GD to GNNs. We observe that perturbed message passing (Equation 2) in GNNs follows a strikingly mathematical parallel pattern as DP-GD: This parallel structure enables to leverage the insight of convergent privacy analysis in the context of GNNs, provided we ensure two critical conditions:
  1. (1)
    Hidden intermediate embeddings: Release only the final node representations ${\bm{X}}^{(K)}$ after $K$ layers, concealing all intermediate states; (Section 4.1)
  2. (2)
    Contractive message passing: Design the message passing operation $\mathrm{MP}_{{\bm{G}}}$ to be provably contractive with coefficient $c<1$, ensuring $\|\mathrm{MP}_{{\bm{G}}}({\bm{X}})-\mathrm{MP}_{{\bm{G}}}({\bm{Y}})\|_{F}\leq c\|{\bm{X}}-{\bm{Y}}\|_{F}$ for all node feature matrices ${\bm{X}},{\bm{Y}}$. (Section 4.2)
When the perturbed message passing step is contractive with respect to the $\ell_{2}$ norm, the distance between GNNs trained on neighboring datasets shrinks at each step. Consequently, the influence of individual data points diminishes, leading to the amplified privacy rooted from “over-smoothing” effect. Accordingly, our insight challenges the previous analysis on private GNNs that accumulates the privacy loss from multi-hop GNN aggregations linearly, and simultaneously removes the over-estimated privacy loss of finally released GNN model to derive a much tighter bound.
4. Private GNNs with Contractiveness
Building on insights in Section 3, we aim to design a message-passing layer that is contractive with a flexibility to adjust the Lipschitz constant, enabling a more efficient and effective privacy budget allocation. To this end, we propose $\mathsf{CARIBOU}$, a modular private GNN framework (Figure 4) with three key modules:
  • 1)
    Contractive Aggregation Module (CAM, § 4.2): We propose a new design of message passing layers with carefully controlled Lipschitz constants to ensure contractiveness so that we can add a reasonable, small yet sufficient noise to protect the computed message passing; Then, we realize a new DP mechanism of perturbed message passing that only releases final node representations ${\bm{X}}^{(K)}$, preventing adversary from exploiting intermediate states and thus amplifying the privacy guarantee rooted in hidden node embeddings.
  • 2)
    Privacy Allocation Module (PAM, § 4.3): PAM ensures efficient privacy budget allocation for edge- and node-level DP guarantees, in which noise calibration is based on new convergent privacy analysis.
  • 3)
    Privacy Auditing Module (PDM, § 4.4): PDM empirically audits graph DP through tests such as membership inference attacks, ensuring theoretical guarantees align with practical deployments.
By integrating these modules, $\mathsf{CARIBOU}$ (Figure 4) achieves accurate and scalable private GNN learning on long-range graphs, with adjustable edge- and node-level privacy under a convergent privacy budget which we will theoretically established in Section 5. Table 6 summarizes acronyms and symbols.
4.1. Clarification on Security Model
Typically, aggregation-based GNNs like GCN (Kipf and Welling, 2017), GCN-II (Chen et al., 2020b), and SAGE (Hamilton et al., 2017) adopt finally obtained node embeddings, while GAP (Sajadmanesh et al., 2023) utilizes the concatenation of all intermediate obtained node embeddings of $K$ layers. This work based on message passing mechanisms adopts the former route to generate node embeddings. Notably, the implicit (also realistic) assumption (Ye and Shokri, 2022) here is that the internal state of $K$-layer message passing is not revealed to the adversary when we design DP algorithms. The final node representation can be applied later in various tasks, including node classification (Wang et al., 2024), link prediction (Rao et al., 2024), or graph classification (Li et al., 2024). Current privacy analysis in private GNNs (e.g., GAP) implies the assumption of revealing all $K$-layer node embeddings computed by perturbed message passing mechanisms to the adversary, which grows the noise scale $\sigma$ linearly ($O(\sqrt{{K}/{\epsilon^{2}}})$) for deriving a theoretical privacy bound. Such privacy bound is loose and we refine it under a more realistic security model (Ye and Shokri, 2022).
CGL Contractive Graph Layer.
RDP Rényi differential privacy.
MIA Membership Inference Attacks.
$K$ The number of layers or hops.
$\epsilon,\sigma$ Privacy parameters: budget, noise scale.
$C_{\mathsf{L}},\Delta$ Lipchitz constant, sensitivity.
$\alpha_{1},\alpha_{2},\beta$ Hyper-parameters of CGL.
Table 1. Acronyms & Symbols
4.2. Contractive Aggregation Module
Contractive operations in GNNs, such as the graph convolution (GConv) layer (Kipf and Welling, [n. d.]), inherently reduce privacy risks by mitigating the memorization of GNN parameters through the over-smoothing phenomenon (Chen et al., 2020a). This property aligns well with the need for $K$-layer aggregation in long-range graphs, where a target node aggregates embeddings from distant source nodes. However, directly stacking $K$ GConv layers, as shown in Figure 3, is suboptimal due to limited expressive power (Chen et al., 2020b) and heightened sensitivity to DP noise (Wu et al., 2019).
To address these limitations, we propose the Contractive Aggregation Module (CAM), centered on the Contractive Graph Layer (CGL). The CGL introduces adjustable coefficients to enhance flexibility and utility while maintaining privacy. Formally, the CGL is defined as:
(4) ${\bm{X}}^{(k+1)}=C_{\mathsf{L}}\big{(}\alpha_{1}\hat{{\bm{A}}}{\bm{X}}^{(k)}+\alpha_{2}\mathsf{Mean}({\bm{X}}^{(k)})\big{)}+\beta{\bm{X}}^{(0)},$
[…]
Analysis of hyperparameters.
The hyperparameter $C_{\mathsf{L}}$ ensures the contractiveness by bounding the magnitude of the output embeddings. Specifically, $C_{\mathsf{L}}$ enforces a Lipschitz constraint, ensuring that small changes in the input embeddings ${\bm{X}}^{(k)}$ result in proportionally small changes in the output ${\bm{X}}^{(k+1)}$. The coefficients $\alpha_{1},\alpha_{2}$ govern the relative contributions of the graph-based aggregation $\hat{{\bm{A}}}{\bm{X}}^{(k)}$ and the mean-based aggregation $\mathsf{Mean}({\bm{X}}^{(k)})$, respectively. By satisfying the constraint $\alpha_{1}+\alpha_{2}=1$, these coefficients ensure a convex combination of the two components. A higher $\alpha_{1}$ emphasizes the influence of the graph topology, leveraging structural information from the adjacency matrix $\hat{{\bm{A}}}$. The parameter $\beta$ controls the residual connection $\beta{\bm{X}}^{(0)}$ calculated independently of the graph topology, incorporating the characteristics of the initial node ${\bm{X}}^{(0)}$ into the output. This residual term mitigates the vanishing gradient problem by preserving a direct path for gradient flow.
Together, the CGL combines three key components: (1) the graph-based aggregation $\hat{{\bm{A}}}{\bm{X}}^{(k)}$, (2) the mean-based aggregation $\mathsf{Mean}({\bm{X}}^{(k)})$, and (3) the residual connection $\beta{\bm{X}}^{(0)}$. This design balances the trade-off between privacy and utility by limiting the propagation of noise while preserving expressive power.
Comparison with existing on general design.
GAP (Sajadmanesh et al., 2023) enhances expressiveness by connecting all intermediate embeddings but lacks contractiveness, compromising utility and privacy. In contrast, CGL introduces adjustable coefficients, achieving a balanced trade-off between privacy, utility, and generalizability. Table 2 summarizes the general design characteristics between GAP and $\mathsf{CARIBOU}$.
4.3. Privacy Allocation Module
For multi-hop aggregation, the features from the previous hop ${\bm{X}}^{(k-1)}$ are aggregated using the adjacency matrix $\hat{{\bm{A}}}$ to enable message passing to neighboring nodes. To ensure privacy, Gaussian noise $\mathcal{N}(\mu,\sigma^{2})$ is added to the aggregated features, where the noise variance $\sigma^{2}$ is determined by the privacy budget $\epsilon$, ensuring compliance with DP guarantees.
Building on Section 4.2, we integrate CAM and PAM to design the Perturbed Contractive Message Passing (PCMP) (Algorithm 1). By leveraging contractiveness, PCMP ensures bounded privacy loss for long $K$-hop graphs, eliminating the linear growth of privacy loss with $K$. A subsequent projection step enforces Lipschitz constraints, maintaining consistent scaling across hops. After $K$ iterations, PCMP outputs private feature matrices $\hat{\mathbf{X}}^{(K)}$, which are passed to the classification module.
Table 2. Comparison between Private Message Passing GNNs. “Utility” summarizes the results of $\mathsf{GAP}$ and $\mathsf{CARIBOU}$ in Table 3. “Chains” is the chain-structured dataset (Gu et al., 2020) and “Standard” represents other real-world graph datasets.
Framework Contractiveness Node Embeddings Calibrated Noise ($\sigma$) Privacy Loss Utility (Chains) Utility (Standard) GAP $\times$ Intermediate & final $\propto\sqrt{K}$ Unbounded $\mathsf{CARIBOU}$checkmark Final $\propto\sqrt{\min(K,\frac{1-C_{\mathsf{L}}^{K}}{1+C_{\mathsf{L}}^{K}}\frac{1+C_{\mathsf{L}}}{1-C_{\mathsf{L}}})}$ Bounded
Algorithm 1 Perturbed Contractive Message Passing (PCMP)
1:Graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with adjacency matrix ${\bm{A}}$; The number of hops $K$; Lipchitz constant $C_{\mathsf{L}}$, DP parameters $\epsilon,\delta$; Initial normalized features $\mathbf{X}^{(0)}$;
2:Private aggregated node embeddings $\hat{{\bm{X}}}^{(K)}$.
3:
4$\triangleright$ Calculate the Required Noise Calibration (from PAM).
5:if Edge-level privacy then
6:     Calculate $\Delta(\mathrm{MP})$ through Equation 6
7:else if Node-level privacy then
8:     Calculate $\Delta(\mathrm{MP})$ through Equation 8
9:end if
10:Calculate $\sigma^{2}$ through Theorem 11
11:
12$\triangleright$ Perturbed Message Passing with Contractiveness (from CAM).
13:for $k=0,\dots,K-1$ do
14:     ${{\bm{X}}}^{(k+1)}\leftarrow C_{\mathsf{L}}(\alpha_{1}\hat{A}{\bm{X}}^{(k)}+\alpha_{2}\mathsf{Mean}({\bm{X}}^{(k)}))+\beta{\bm{X}}^{(0)}$
15:     $\triangleright$ Contractive graph layer: compute the node embeddings.
16:     ${{\bm{X}}}^{(k+1)}\leftarrow{{\bm{X}}}^{(k+1)}+\mathcal{N}(\mu,(\Delta(\mathsf{CGL}))^{2}\sigma^{2})$triangleright Perturbation.
17:     ${{\bm{X}}}^{(k+1)}\leftarrow\Pi_{\mathcal{K}}({{\bm{X}}}^{(k)})$triangleright Projection constrained by norm $1$.
18:end for
19:
20:Return: ${\mathbf{X}}^{(K)}$
Privacy Budget Allocation.
The privacy budget is distributed across $K$ hops to ensure the total privacy loss adheres to the specified $\epsilon$ and $\delta$. The noise scale is calibrated based on the sensitivity of graph operations and the desired DP guarantees, using a noise allocation mechanism (NAM) that limits noise accumulation under Lipschitz constraints. The maximum allowable noise calibration (see Corollary 10) is constrained by $K^{\prime}=\min(K,(1+C_{\mathsf{L}})/(1-C_{\mathsf{L}}))$.
PCMP integrates contractive aggregation and privacy-preserving perturbation for private message passing. Noise calibration begins by determining the sensitivity of graph operations and computing the noise scale. At each hop, the CGL aggregates features while maintaining contractiveness through adjustable coefficients. Gaussian noise is added to ensure privacy, and embeddings are normalized to enforce Lipschitz constraints. This iterative process produces private embeddings suitable for downstream tasks.
4.4. Privacy Auditing Module
Message passing mechanisms integrate graph structures by recursively aggregating and transforming node features from neighbors. Membership Inference Attacks (MIA) on graph examine the vulnerability of message passing to infer whether specific nodes or edges were part of a GNN’s training set (Olatunji et al., 2021b). The adversaries exploit black-box access to the GNN, querying it with selected data and analyzing outputs (e.g., class probabilities) to infer membership. To align theoretical privacy guarantees with practical deployments, we propose an empirical auditing module. This module simulates an adversary to evaluate GNN privacy before deployment. Extending Carlini et al. (Carlini et al., 2022)’s MIA framework to graphs, we define a graph-specific privacy auditing game (Definition 1) and implement two real-world attacks (Wu et al., 2022a; Olatunji et al., 2021b) for privacy verification.
Definition 0 (Graph-based Privacy Auditing Game).
The game proceeds between a challenger $\mathcal{C}$ and an privacy auditor $\mathcal{A}$:
  1. (1)
    The challenger samples a training graph in the transductive setting (a set of subgraphs in the inductive setting) $G\leftarrow\mathbb{G}$ and trains a model $f_{\theta}\leftarrow\mathcal{T}(G)$ on the dataset $G$.
  2. (2)
    The challenger flips a bit $b$, and if $b=0$, samples a fresh challenge point from the distribution $(x,y)\leftarrow\mathbb{G}$ (such that $(x,y)\notin G$). Otherwise, the challenger selects a point from the training set $(x,y)\overset{\$}{\leftarrow}G$.
  3. (3)
    The challenger sends $(x,y)$ to the adversary.
  4. (4)
    The adversary gets query access to the distribution $\mathbb{G}$, and to the model $f_{\theta}$, and outputs a bit $\hat{b}\leftarrow\mathcal{A}^{\mathbb{G},f}(x,y)$.
  5. (5)
    Output 1 if $\hat{b}=b$, and 0 otherwise.
The attacker can output either a “hard prediction,” or a continuous confidence score, thresholded as a reference to yield a membership prediction.
4.5. Putting it Together
As shown in Figure 4, before and after the perturbed message passing, $\mathsf{CARIBOU}$ employs an encoder and a classification module (CM) to process the node features. To ensure compliance with DP guarantees, the framework utilizes standard DP-SGD during pre-training.
Upon completing the private $K$-hop aggregations, the resulting private graph embeddings are passed to the CM. The CM integrates two key components: (1) the graph-agnostic node features ${\bm{X}}^{(0)}$, which capture individual node characteristics independent of the graph structure, and (2) the private, topology-aware aggregated features ${\bm{X}}^{\prime(K)}$, which encode structural information from the graph. This dual integration enhances the model’s expressiveness while preserving privacy.
To improve classification accuracy, $\mathsf{CARIBOU}$ adopts a head MLP architecture proposed by Sajadmanesh et al. (Sajadmanesh et al., 2023) as the CM. This design ensures that the CM effectively combines the information from both feature sets, enabling robust node classification. Furthermore, the CM guarantees efficient training by leveraging the graph-agnostic features ${\bm{X}}^{(0)}$, ensuring a lower-bound performance even in scenarios where graph topology is unavailable.
5. Convergent Privacy Analysis
In this section, we present a convergent privacy analysis for perturbed message-passing GNNs with respect to the number of hops. We begin by reviewing the standard privacy analysis for a one-layer perturbed message-passing GNN, and then illustrate that the privacy cost grows linearly with the number of layers under standard composition theorems. We then shift our focus to a convergent privacy analysis for perturbed GNNs whose message-passing layers are contractive. In particular, we draw upon the framework of contractive noisy iteration (CNI) from Bok et al. (2024), recasting the multi-layer perturbed GNN as a CNI process. Our analysis reveals that, under hidden intermediate states and contractive message-passing layers, the privacy cost converges as the number of hops increases. Finally, we specialize this result to our proposed $\mathsf{CARIBOU}$. We show that $\mathsf{CARIBOU}$’s message-passing operation is contractive, derive its sensitivity, and thereby establish concrete bounds on both edge-level and node-level differential privacy for arbitrarily many hops. All proofs are deferred to the Appendix C.
5.1. Preliminaries: Perturbed Message Passing
Recall that each layer of a perturbed message-passing GNN updates the node features following Equation 2, where ${\bm{Z}}\sim\mathcal{N}\bigl{(}0,\sigma^{2}\mathrm{I}_{d}\bigr{)}$ is Gaussian noise added to obscure the true output of $\mathrm{MP}_{{\bm{G}}},$ and $\Pi_{\mathcal{K}}$ projects onto some convex feasible set ${\mathcal{K}}.$ As shown in Proposition 1, for a single layer, the privacy cost depends only on $\Delta(\mathrm{MP})$ and $\sigma.$ However, by the composition theorem of RDP (Mironov, 2017), stacking $K$ such layers yields a cost that scales linearly in $K$ (Corollary 2).
Proposition 0.
[…]
Corollary 0.
[…]
This linear growth of the privacy budget is common in multi-hop settings, such as (Sajadmanesh et al., 2023), but is troublesome for deep GNNs, especially when many message-passing steps are needed for long-range dependencies. Excessively large privacy loss often undermines the model’s utility.
5.2. Contractive Noisy Iteration and Convergent Privacy
Many GNNs, such as GCN, use "mean-type" aggregation, mixing a node’s representation with that of its neighbors. Intuitively, iterative mixing could "amplify" privacy, but existing analyses yield only linear compositions. Our key observation is that perturbed message passing in a contractive GNN layer behaves analogously to noisy gradient descent or noisy iterative maps (Chourasia et al., 2021; Altschuler and Talwar, 2022; Altschuler et al., 2024), where recent work has demonstrated privacy amplification via iteration.
Below, we introduce the framework of contractive noisy iteration (CNI) from Bok et al. (2024), and the meta theorem proved by Bok et al. (2024) that provides a tight privacy guarantee for CNI processes.
Definition 0 (Contractive Noisy Iteration (CNI)(Bok et al., 2024, Definition 3.1)).
Consider a sequence of random variables
(5) ${\bm{X}}^{(k+1)}=\Pi_{\mathcal{K}}\!\bigl{(}\,\phi_{k+1}\bigl{(}{\bm{X}}^{(k)}\bigr{)}+{\bm{Z}}^{(k)}\bigr{)},$
where each map $\phi_{k}$ is Lipchitz continuous, ${\bm{Z}}^{(k)}\sim\mathcal{N}\bigl{(}0,\,\sigma^{2}\,\mathrm{I}\bigr{)}$ are i.i.d. Gaussian noise vectors independent of ${\bm{X}}^{(k)}$, and $\Pi_{\mathcal{K}}$ is a projection operator onto a convex feasible set $\mathcal{K}\subseteq\mathbb{R}^{d}.$ This iterative process is called contractive noisy iteration (CNI).
A special case of CNI considered in (Bok et al., 2024) is the noisy gradient descent, where the contractive functions are the gradient update steps for some fixed loss function $f$, and the noise distribution $\xi_{k}$ is Gaussian. This situation is similar to the perturbed message passing mechanism in a contractive GNN layer, where the contractive function $\phi_{k}\equiv\mathrm{MP}_{G}$ is the message passing operation for a fixed graph ${\bm{G}}$.
The complete privacy analysis of CNI processes in Bok et al. (2024) involves concepts of trade-off functions (Definition 1) and Gaussian differential privacy (GDP, Definition 2), which we briefly review in Appendix B. GDP can be converted to the more familiar Rényi Differential Privacy (RDP) framework:
Lemma 0 (GDP Implies RDP (Bok et al., 2024, Lemma A.4)).
[…]
With this connection established, we now state the key meta theorem from Bok et al. (2024) that analyzes the CNI process and provides a tight privacy guarantee.
Theorem 5 (Meta Theorem on CNI (Bok et al., 2024, Theorem C.5)).
Let $\{{\bm{X}}^{(k)}\}_{k=0}^{K}$ and $\{{\bm{X}}^{\prime(k)}\}_{k=0}^{K}$ represent the output of CNI processes,
$\text{CNI}({\bm{X}}^{(0)},\{\phi_{k}\}_{k=1}^{K},\{\mathcal{N}(0,\sigma^{2}{\bm{I}}_{d})\}_{k=1}^{K},{\mathcal{K}}),\text{ and}$
$\text{CNI}({\bm{X}}^{(0)},\{\phi_{k}^{\prime}\}_{k=1}^{K},\{\mathcal{N}(0,\sigma^{2}{\bm{I}}_{d})\}_{k=1}^{K},{\mathcal{K}}).$
Assume that:
  • $\phi_{1}$ and $\phi^{\prime}_{1}$ are Lipschitz continuous,
  • each $\phi_{k},\phi_{k}^{\prime}$ is $\gamma$-Lipschitz, with $\gamma<1$ for $k=2,\dots,K$,
  • $\|\phi_{k}(x)-\phi_{k}^{\prime}(x)\|\leq s$ for all $x$ and $k=1,\dots,K$,
Then the tradeoff function $T({\bm{X}}^{(K)},{\bm{X}}^{\prime(K)})$ satisfies,
$T({\bm{X}}^{(K)},{\bm{X}}^{\prime(K)})\geq T({\mathcal{N}}(0,1),{\mathcal{N}}(\mu^{(K)},1)),$
where,
$\mu^{(K)}=\sqrt{\frac{1-\gamma^{K}}{1+\gamma^{K}}\frac{1+\gamma}{1-\gamma}}\frac{s}{\sigma}.$
Remark 1.
The theorem above slightly generalizes the original result from Bok et al. (2024) by relaxing the Lipschitz condition to require $\gamma<1$ only for $k\geq 2$ rather than for all iterations. This relaxation is critical for analyzing our $\mathsf{CARIBOU}$ architecture, where the first message-passing layer includes the residual term $\beta{\bm{X}}^{(0)}$, potentially making it non-contractive while subsequent layers remain contractive. The proof extends the original argument by carefully tracking the influence of the first layer on the privacy guarantee. For completeness, we provide the full proof in Appendix C.
By utilize the above meta theorem and property of Gaussian tradeoff function, we can derive the privacy guarantee for a $K$-layer perturbed message-passing GNN with contractive message passing layers.
Theorem 6 ((Bok et al., 2024, Theorem 4.2) adapted for GNNs).
Let $\mathrm{MP}$ be a message passing mechanism of a GNN such that the message passing operator $\mathrm{MP}_{G}$ for any graph ${\bm{G}}$ is contractive with Lipschitz constant $\gamma<1$ for layers $k\geq 2$. Assume the sensitivity of $\mathrm{MP}$ is $\Delta(\mathrm{MP})$ and the noise scale is $\sigma$. Then, a $K$-layer perturbed message passing GNN with $\mathrm{MP}$ satisfies,
[…]
The upshot is that as $K\to\infty$, the privacy parameter converges to $\frac{1}{2}\alpha\frac{\Delta^{2}(\mathrm{MP})}{\sigma^{2}}\frac{1+\gamma}{1-\gamma}$, rather than growing unbounded with $K$ as in standard composition. This result enables meaningful privacy guarantees even for deep GNNs with many message-passing layers.
5.3. Edge- and Node-Level Privacy of $\mathsf{CARIBOU}$
We now specialize to the $\mathsf{CARIBOU}$ architecture and establish both edge-level and node-level DP guarantees. Recall the contractive graph layer (CGL) of $\mathsf{CARIBOU}$:
[…]
where $10\leq C_{\mathsf{L}}<$ 1, and hyper-parameters, $\alpha_{1},\alpha_{2},\beta>0,\alpha_{1}+\alpha_{2}=1$. In order to establish the privacy guarantees of $\mathsf{CARIBOU}$, we need to determine the sensitivity and the Lipschitz constant of the message passing layer $\mathsf{CGL}$.
Proposition 0.
The message passing layer $\mathsf{CGL}$ of $\mathsf{CARIBOU}$ is contractive with Lipschitz constant $C_{\mathsf{L}}<1$ with respect to the input ${\bm{X}}^{(k)}$, for any $k\geq 2$.
The sensitivity of the message passing layer $\mathsf{CGL}$ regarding edge-level and node-level adjacency graphs is determined as follows.
Theorem 8 (Edge-level Sensitivity of $\mathsf{CGL}$).
Let ${\bm{G}}$ be a graph and $D_{\min}$ be the minimum node degree of ${\bm{G}}$. The edge level sensitivity $\Delta_{e}$ of the message passing layer $\mathsf{CGL}$ is
(6) $\Delta_{e}(\mathsf{CGL})=\sqrt{2}C_{\mathsf{L}}\alpha_{1}\Big{(}$ $\frac{1}{(D_{\min}+1)(D_{\min}+2)}$
$+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}\Big{)},$
where $C(D_{\min})$ is a piecewise function defined as
(7) $C(D_{\min})=\begin{cases}\frac{D_{\min}}{\sqrt{D_{\min}+1}}-\frac{D_{\min}}{\sqrt{D_{\min}+2}}&D_{\min}>3\\ (\frac{3}{\sqrt{4}}-\frac{3}{\sqrt{5}})&1\leq D_{\min}\leq 3\\ \end{cases}$
Intuitively, the edge-level sensitivity $\Delta_{e}$ will be smaller if the minimum degree $D_{\min}$ is larger or the Lipschitz constant $C_{\mathsf{L}}$ is smaller. The node level sensitivity $\Delta_{n}$ of the message passing layer $\mathsf{CGL}$ is determined as follows.
Theorem 9 (Node-level Sensitivity of $\mathsf{CGL}$).
Let ${\bm{G}}$ be a graph and $D_{\max}$ be the maximum node degree of ${\bm{G}}$. The node level sensitivity $\Delta_{n}$ of the message passing layer $\mathsf{CGL}$ is
(8) $\Delta_{n}(\mathsf{CGL})=1+\quad\alpha_{2}C_{\mathsf{L}}\frac{2|V|}{|V|+1}$
$+\alpha_{1}C_{\mathsf{L}}\left(\frac{\sqrt{D_{\max}}}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})\sqrt{D_{\max}}}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}}\right)$
where $C(D_{\min})$ is defined as in equation 7.
From the above, we can see that the node-level sensitivity $\Delta_{n}$ will be smaller if the maximum degree $D_{\max}$ is smaller, the minimum degree $D_{\min}$ is larger, or the Lipschitz constant $C_{\mathsf{L}}$ is smaller.
Then as a consequence of the contractive result in Proposition 7 and the sensitivity results in Theorem 8 and Theorem 9, we can apply Theorem 6 to derive the following privacy guarantee for the message-passing layer $\mathsf{CGL}$ of $\mathsf{CARIBOU}$.
Corollary 0 (DP Guarantees for $\mathsf{CGL}$ layers).
Let ${\bm{G}}$ be a graph and $K$ be the number of hops ($\mathsf{CGL}$ layers) in $\mathsf{CARIBOU}$. Then the $K$-hop message passing of $\mathsf{CARIBOU}$ satisfies:
[…]
For a complete analysis, we integrate the node-level privacy guarantees into an overall DP bound for the entire training process of $\mathsf{CARIBOU}$. Specifically, we assume that the private DAE and CM satisfy $(\alpha,\epsilon_{\mathsf{DAE}}(\alpha))$-RDP and $(\alpha,\epsilon_{\mathsf{CM}}(\alpha))$-RDP, respectively. By privacy composition, $\mathsf{CARIBOU}$’s overall privacy budget $\epsilon_{\mathsf{CARIBOU}}$ is then derived as shown in Theorem 11. The terms $\epsilon_{\mathsf{DAE}}(\alpha)$ and $\epsilon_{\mathsf{CM}}(\alpha)$ individually quantify the privacy contributions from the DAE and CM modules, while the remaining aggregation term accounts for node-level privacy loss during $K$-hop message passing. Finally, the $\frac{\log(1/\delta)}{\alpha-1}$ term incorporates the privacy failure probability $\delta$, ensuring conventional privacy guarantees across the inference process.
Theorem 11 ($\mathsf{CARIBOU}$’s Privacy Guarantee).
For any $\alpha>1$, let DAE training and CM training satisfy $(\alpha,\epsilon_{\mathsf{DAE}}(\alpha))$-RDP and $(\alpha,\epsilon_{\mathsf{CM}}(\alpha))$-RDP, respectively. Then, for the maximum hop $K\geq 1$, privacy failure probability $10<\delta<$ 1, Lipschitz constant $10<C_{\mathsf{L}}<$ 1, and noise variance $\sigma^{2}$, $\mathsf{CARIBOU}$ satisfies $(\epsilon_{\mathsf{CARIBOU}},\delta)$-DP, where $\epsilon_{\mathsf{CARIBOU}}=\epsilon_{\mathsf{DAE}}(\alpha)+\epsilon_{\mathsf{CM}}(\alpha)+\frac{\alpha}{2}\frac{\Delta^{2}}{\sigma^{2}}\min\left\{K,\frac{1-C_{\mathsf{L}}^{K}}{1+C_{\mathsf{L}}^{K}}\frac{1+C_{\mathsf{L}}}{1-C_{\mathsf{L}}}\right\}+\frac{\log(1/\delta)}{\alpha-1}.$ Here, $\Delta=\Delta_{e}(\mathsf{CGL})$ from Theorem 8 for edge-level privacy, or $\Delta=\Delta_{n}(\mathsf{CGL})$ from Theorem 9 for node-level privacy.
Table 3. Comparison of classification accuracy. The best accuracy and the second-best accuracy are highlighted, respectively. The symbol
[…]
Dataset Chain-S Chain-M Chain-L Chain-X Computers Facebook PubMed Cora Photo $\epsilon=1$ $\mathsf{CARIBOU}$ $84.4\%$  $82.5\%$  $70.0\%\mathrel{\hskip 5.12128pt}$ $66.0\%\mathrel{\hskip 5.12128pt}$ $92.0\%$ $74.0\%\mathrel{\hskip 5.12128pt}$ $88.2\%$ $85.1\%\mathrel{\hskip 5.12128pt}$ $95.8\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.0\%$ $60.6\%\mathrel{\hskip 5.12128pt}$ $88.3\%$ $75.5\%\mathrel{\hskip 5.12128pt}$ $92.5\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $67.5\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $67.0\%\mathrel{\hskip 5.12128pt}$ $87.2\%$ $68.5\%\mathrel{\hskip 5.12128pt}$ $87.4\%$ $77.1\%\mathrel{\hskip 5.12128pt}$ $93.0\%$ SAGE $53.1\%$blacktriangledown $45.0\%$blacktriangledown $53.3\%$blacktriangledown $51.0\%$blacktriangledown $77.8\%$ $48.2\%$blacktriangledown $85.0\%$ $60.0\%\mathrel{\hskip 5.12128pt}$ $82.4\%$ $\epsilon=2$ $\mathsf{CARIBOU}$ $90.6\%$ $82.5\%$ $73.3\%$ $68.0\%\mathrel{\hskip 5.12128pt}$ $92.1\%$ $73.8\%\mathrel{\hskip 5.12128pt}$ $89.1\%$ $86.9\%\mathrel{\hskip 5.12128pt}$ $95.9\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.2\%$ $66.7\%\mathrel{\hskip 5.12128pt}$ $88.2\%$ $77.5\%\mathrel{\hskip 5.12128pt}$ $93.3\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $67.5\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $68.0\%\mathrel{\hskip 5.12128pt}$ $88.3\%$ $71.7\%\mathrel{\hskip 5.12128pt}$ $87.5\%$ $77.5\%\mathrel{\hskip 5.12128pt}$ $93.4\%$ SAGE $43.8\%$blacktriangledown $45.0\%$blacktriangledown $53.3\%$blacktriangledown $51.0\%$blacktriangledown $76.1\%$ $48.1\%$blacktriangledown $84.6\%$ $60.1\%\mathrel{\hskip 5.12128pt}$ $82.4\%$ $\epsilon=4$ $\mathsf{CARIBOU}$ $90.6\%$ $82.5\%$ $73.3\%$ $69.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $74.0\%\mathrel{\hskip 5.12128pt}$ $89.5\%$ $87.3\%$ $95.9\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.9\%$ $73.3\%\mathrel{\hskip 5.12128pt}$ $88.4\%$ $75.1\%\mathrel{\hskip 5.12128pt}$ $94.2\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $67.5\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $67.0\%\mathrel{\hskip 5.12128pt}$ $87.2\%$ $68.5\%\mathrel{\hskip 5.12128pt}$ $87.4\%$ $77.1\%\mathrel{\hskip 5.12128pt}$ $93.0\%$ SAGE $50.0\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $54.0\%$blacktriangledown $79.1\%$ $50.3\%$blacktriangledown $85.8\%$ $63.3\%\mathrel{\hskip 5.12128pt}$ $85.7\%$ $\epsilon=8$ $\mathsf{CARIBOU}$ $93.8\%$ $82.5\%$ $75.0\%$ $70.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $74.4\%\mathrel{\hskip 5.12128pt}$ $89.8\%$ $88.4\%\mathrel{\hskip 5.12128pt}$ $96.0\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $89.4\%$ $78.6\%\mathrel{\hskip 5.12128pt}$ $88.6\%$ $76.0\%\mathrel{\hskip 5.12128pt}$ $94.6\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $60.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $67.0\%\mathrel{\hskip 5.12128pt}$ $90.1\%$ $75.0\%\mathrel{\hskip 5.12128pt}$ $88.1\%$ $79.0\%\mathrel{\hskip 5.12128pt}$ $94.6\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $61.0\%\mathrel{\hskip 5.12128pt}$ $87.9\%$ $75.6\%\mathrel{\hskip 5.12128pt}$ $84.8\%$ $75.7\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $\epsilon=16$ $\mathsf{CARIBOU}$ $93.8\%$ $85.0\%$ $73.3\%$ $73.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $74.5\%\mathrel{\hskip 5.12128pt}$ $89.8\%$ $88.6\%\mathrel{\hskip 5.12128pt}$ $96.0\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $40.0\%$blacktriangledown $90.4\%$ $81.2\%\mathrel{\hskip 5.12128pt}$ $88.9\%$ $77.7\%\mathrel{\hskip 5.12128pt}$ $93.8\%$ GAP $62.5\%\mathrel{\hskip 5.12128pt}$ $70.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $68.0\%\mathrel{\hskip 5.12128pt}$ $90.1\%$ $76.0\%\mathrel{\hskip 5.12128pt}$ $88.5\%$ $81.7\%\mathrel{\hskip 5.12128pt}$ $94.6\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $51.0\%$blacktriangledown $90.9\%$ $79.5\%\mathrel{\hskip 5.12128pt}$ $87.6\%$ $84.7\%\mathrel{\hskip 5.12128pt}$ $94.1\%$ $\epsilon=32$ $\mathsf{CARIBOU}$ $93.8\%$ $82.5\%$ $78.3\%$ $73.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $74.0\%\mathrel{\hskip 5.12128pt}$ $89.9\%$ $88.6\%\mathrel{\hskip 5.12128pt}$ $95.9\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $40.0\%$blacktriangledown $91.3\%$ $82.9\%\mathrel{\hskip 5.12128pt}$ $88.8\%$ $79.9\%\mathrel{\hskip 5.12128pt}$ $94.3\%$ GAP $59.4\%\mathrel{\hskip 5.12128pt}$ $65.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $68.0\%\mathrel{\hskip 5.12128pt}$ $90.6\%$ $77.0\%\mathrel{\hskip 5.12128pt}$ $88.8\%$ $82.5\%\mathrel{\hskip 5.12128pt}$ $94.8\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $51.0\%$blacktriangledown $90.6\%$ $79.9\%\mathrel{\hskip 5.12128pt}$ $86.9\%$ $85.2\%\mathrel{\hskip 5.12128pt}$ $94.4\%$ $\mathsf{CARIBOU}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $92.4\%$ $79.0\%\mathrel{\hskip 5.12128pt}$ $89.9\%$ $89.3\%\mathrel{\hskip 5.12128pt}$ $96.0\%$ DPDGC $59.4\%\mathrel{\hskip 5.12128pt}$ $77.5\%\mathrel{\hskip 5.12128pt}$ $63.3\%\mathrel{\hskip 5.12128pt}$ $73.0\%\mathrel{\hskip 5.12128pt}$ $92.8\%$ $86.4\%\mathrel{\hskip 5.12128pt}$ $88.1\%$ $83.9\%\mathrel{\hskip 5.12128pt}$ $96.2\%$ Plain GAP $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $91.0\%$ $79.0\%\mathrel{\hskip 5.12128pt}$ $89.3\%$ $85.2\%\mathrel{\hskip 5.12128pt}$ $95.5\%$ $\epsilon=\infty$ SAGE $59.4\%\mathrel{\hskip 5.12128pt}$ $55.0\%$blacktriangledown $60.0\%\mathrel{\hskip 5.12128pt}$ $58.0\%\mathrel{\hskip 5.12128pt}$ $91.6\%$ $79.7\%\mathrel{\hskip 5.12128pt}$ $87.0\%$ $82.1\%\mathrel{\hskip 5.12128pt}$ $94.2\%$ MLP $40.6\%$blacktriangledown $47.5\%$blacktriangledown $46.7\%$blacktriangledown $51.0\%$blacktriangledown $83.9\%$ $51.1\%$blacktriangledown $84.9\%$ $72.9\%\mathrel{\hskip 5.12128pt}$ $91.1\%$
6. Experimental Evaluation
The empirical evaluation includes privacy-utility trade-offs, privacy audits, ablation studies of hops and hyper-parameters, and computational overhead.
Datasets.
$\mathsf{CARIBOU}$ was tested over nine graph datasets. Five of the datasets have been broadly used to evaluate GNN methods, including Photo and Computers (Shchur et al., 2018), Cora and PubMed (Sen et al., 2008), Facebook (Traud et al., 2011). We also adjusted the synthetic chain-structured dataset developed under (Gu et al., 2020) into various scales, termed Chain-S, Chain-M, Chain-L and Chain-X. The chain-structured dataset has been used to understand the relations between privacy/utility and hops, as described in Section 3.1. It is considered as an important benchmark to assist the development of long-range interaction GNNs by the ML community (Dwivedi et al., 2022). More details on datasets are specialized in Appendix D.1.
Baselines.
We compare $\mathsf{CARIBOU}$ with multiple edge-level private GNNs and vanilla GNNs (Chien et al., 2023; Sajadmanesh et al., 2023; Wu et al., 2022b; Daigavane et al., 2021). In particular, we use GAP (Sajadmanesh et al., 2023) and DPDGC (Chien et al., 2023) as the baseline private GNN methods, as both have open-source implementations. SAGE and MLPare non-private baseline methods and we adapt them to the private settings. More details of the baseline methods are provided in Appendix D.2.
Model Architectures & Configurations of $\mathsf{CARIBOU}$.
We concatenate $K$ layers ranging from $1$ to $20$, with $K,\alpha_{1},\beta,C_{\mathsf{L}}$ as the hyper-parameters. If the range between two interactive nodes is long, we use larger values to test the influence of hops. Specifically, for Chain-S  and Chain-M  datasets, $\mathsf{CARIBOU}$ uses the number of hops $K\in\{15,10,9,8,7,6,5,4,3,2,1\}$, while for Chain-S  and Chain-M  datasets consisting of longer chains, $\mathsf{CARIBOU}$ takes $K\in\{20,17,15,\allowbreak 13,11,9,7,5\}$. As for standard datasets listed in Table 5, $K$ is chosen from the union sets of the aforementioned two sets, i.e., $\{20,17,15,\allowbreak 13,11,10,9,8,7,6,5,4,3,2,1\}$. To align with GAP’s configuration (Sajadmanesh et al., 2023), we set the number of hidden units to $\{16,64\}$ and use the SeLU activation function (Klambauer et al., 2017) at every layer. We adopt the Adam optimizer over $100$ epochs with a learning rate $0.001$, and pick the best accuracy to report.
Privacy Configuration and Parameters.
We implement convergent privacy allocation based on GAP’s privacy budget accounting mechanism (Sajadmanesh et al., 2023) and numerically calibrate noise level $\sigma$ by setting $\epsilon$. All edge/node-level private algorithms adopt the Gaussian mechanism to sample the i.i.d. noise to realize the desired $(\epsilon,\delta)$-DP. For comprehensive evaluation, we consider different choices of $\epsilon\in\{1,2,4,8,16,32\}$, while $\delta$ is set to be smaller than the inverse number of edges for EDP or the inverse number of node for NDP.
Hardware and software.
Our experiments have been conducted on the Ubuntu 20.04.2 LTS server, with AMD Ryzen Threadripper 3970X 32-core CPUs of 256 GB CPU memory and NVIDIA GeForce RTX 3090 of 24GB memory. $\mathsf{CARIBOU}$ equipped with edge- and node-level privacy has been implemented in PyTorch using PyTorch-Geometric (PyG) (Fey and Lenssen, 2019) framework. DP implementation adopts the autodp library (Wang, 2025), which includes analytical moments accountant and private training with DP-SGD (Abadi et al., 2016).
Table 4. Privacy Auditing via Linkteller and G-MIA. AUC score is reported.
Dataset Linkteller G-MIA ϵ=infitalic-$\epsilon=\inf$ $\epsilon=1$ $\epsilon=2$ $\epsilon=4$ $\epsilon=8$ $\epsilon=16$ $\epsilon=32$ ϵ=infitalic-$\epsilon=\inf$ $\epsilon=1$ $\epsilon=2$ $\epsilon=4$ $\epsilon=8$ $\epsilon=16$ $\epsilon=32$ Facebook $0.977$ $0.463$ $0.483$ $0.482$ $0.478$ $0.472$ $0.462$ $0.567$ $0.587$ $0.587$ $0.587$ $0.583$ $0.587$ $0.588$ PubMed $0.981$ $0.446$ $0.442$ $0.441$ $0.445$ $0.442$ $0.449$ $0.600$ $0.599$ $0.598$ $0.598$ $0.601$ $0.601$ $0.605$ Cora $0.998$ $0.427$ $0.448$ $0.399$ $0.443$ $0.451$ $0.450$ $0.645$ $0.500$ $0.500$ $0.500$ $0.500$ $0.500$ $0.500$ Photo $0.962$ $0.475$ $0.404$ $0.421$ $0.428$ $0.417$ $0.434$ $0.678$ $0.677$ $0.682$ $0.682$ $0.678$ $0.672$ $0.676$ Computers $0.860$ $0.367$ $0.364$ $0.372$ $0.363$ $0.361$ $0.384$ $0.702$ $0.701$ $0.661$ $0.707$ $0.708$ $0.711$ $0.701$
Computational overhead.
We elaborate the overhead analysis, including latency and memory, in Appendix E.1.
6.1. Trade-offs of Privacy and Accuracy
We first compare $\mathsf{CARIBOU}$ with the baseline methods on all datasets for their downstream classification tasks. We focus on EDP here as this is our primary privacy setting and leave the results of NDP in Appendix E.2. We run each model $3$ times for each group of hyper-parameters and report the top classification accuracy in Table 3. The mean accuracy over the 3 runs is shown in Table 9 of Appendix.
First, we highlight the results on the chain-structured datasets. In summary, we observe that the non-private versions (i.e., ϵ=infitalic-$\epsilon=\inf$) of GAP and $\mathsf{CARIBOU}$ can achieve perfect classification accuracy, while DPDGC, SAGE and MLP show relatively low accuracy. The root cause is that features of a distant node can only be learned through a sufficient number of hops of message passing, i.e., aggregating and passing features of the neighboring node one by one, but DPDGC and SAGE are not designed to leverage long-hop interactions. For MLP, as the graph structure is not utilized at all, the result is close to random guessing. After adding DP noises to protect edge-level memberships, the accuracy of both GAP and $\mathsf{CARIBOU}$ drop, but the accuracy loss of GAP is more prominent (from $100\%$ down to $57.8\%$-$62.5\%$). GAP is able to maintain reasonable accuracy even with fairly tight privacy budget $\epsilon=1$ on small chain-structured datasets like Chain-S and Chain-M. On the other hand, Chain-L and Chain-X have relatively low accuracy even for $\mathsf{CARIBOU}$ (e.g., $70.0\%$ and $66.0\%$ when $\epsilon=1$). This observed phenomenon is attributed to the increased difficulty of GNN classification with longer chains, as DP noise is injected at every layer, compounding its impact.
Regarding the other standard graph datasets, for Computers, PubMed, Cora and Photo, $\mathsf{CARIBOU}$ can outperform all the other baselines in most cases with varying $\epsilon$. It is also worthy noting that $\mathsf{CARIBOU}$ achieves fairly stable accuracy with with varying $\epsilon$ (e.g., $92.0\%$ to $92.4\%$ for Computers with $\epsilon$ growing from $1$ to $\inf$), suggesting the convergent privacy design is indeed effective. For Facebook, we found $\mathsf{CARIBOU}$ outperforms the other baselines with small $\epsilon$ (i.e., $\epsilon\leq 4$), but falls behind at larger $\epsilon$. Yet, we argue that performance at low privacy budget is more critical for privacy-sensitive datasets like Facebook, and stable accuracy across privacy budget $\epsilon$ might be more desired.
Finding 1.
While GAP demonstrates advantages over existing baselines across various chain-structured datasets and parameter regimes, $\mathsf{CARIBOU}$ further advances GAP’s performance, achieving a more favorable privacy-utility trade-off.
6.2. Privacy Auditing
Following the framework outlined in Section 4.4, we implement two attack methods, Linkteller (Wu et al., 2022a) and G-MIA (Olatunji et al., 2021b), for privacy auditing. Linkteller focuses on edge-level privacy in a training-then-inference scenario. During the inference phase, Linkteller infers the private edge information held by graph holder via influence analysis and exploits adversarial queries from the feature holder. We adopt G-MIA’s attacking settings of TSTF, where models have been trained on subgraph and tested on full graph. The adversary knows the whole graph ${\bm{G}}$ and all edges contained in ${\bm{G}}$ but has no access to the subgraphs used for early training.
In Table 4, we report the AUC (Area Under the Curve) score about the attack effectiveness, when $\mathsf{CARIBOU}$ is being attacked. AUC is a major metric to evaluate the membership inference attack (Wu et al., 2022a; He et al., 2021a). Specifically, AUC measures true positive rate against the false positive rate on various classification thresholds, and a score of $0.5$ suggesting random guessing. We found $\mathsf{CARIBOU}$ is very effective against Linkteller, by dropping the attack AUC from between 0.86 to 0.998 across all standard datasets (ϵ=infitalic-$\epsilon=\inf$) to less than 0.500 ($\epsilon$ ranges from 1 to 32). For G-MIA, its AUC on the 5 datasets are already lower than Linkteller by a notable margin when ϵ=infitalic-$\epsilon=\inf$ (between 0.567 to 0.702 for the 5 datasets), so the impact of $\mathsf{CARIBOU}$ is relative small. But we observe on Cora, $\mathsf{CARIBOU}$ is able to drop AUC from 0.645 to 0.500.
Finding 2.
In privacy auditing, $\mathsf{CARIBOU}$’s shows effective resistance to membership inference attacks.
6.3. Ablation Study
Impact of $K$.
Both $\mathsf{CARIBOU}$ and GAP perform $K$-hop aggregations under $K$ aggregation layers. Here we evaluate the impact of $K$ on accuracy on the chain-structured datasets (Chain-S, Chain-M and Chain-L), as their classification results highly depend on long-range interactions. In Figure 8 of Appendix, we compare $\mathsf{CARIBOU}$ and GAP on varying $\epsilon$. The result of GAP is drawn with lines and the result of $\mathsf{CARIBOU}$ is illustrated with the colored boxes, because $\mathsf{CARIBOU}$ also depends on other hyper-parameters $C_{\mathsf{L}},\alpha_{1},\beta$ and we use the colored boxes to represent the interquartile range over 5 runs of their different value combinations.
In Figure 5(a), we show the result of one setting ($\epsilon=1$ on Chain-S), and $\mathsf{CARIBOU}$ achieves higher accuracy at every $K$. This observation is consistent in other settings as shown in Figure 8. On the other hand, the classification accuracy is fluctuated when $K$ increase, for both $\mathsf{CARIBOU}$ and GAP. For instance, the best accuracy of $\mathsf{CARIBOU}$ is achieved when $K$ equals to $6$ or $7$, which is close to the chain length (i.e., $8$ for Chain-S). This result suggests the added DP noises could outweigh the benefits of aggregations at some point, though $\mathsf{CARIBOU}$ handles this issue better than GAP.
Figure 5. Comparison between $\mathsf{CARIBOU}$ (colored boxes) and GAP (blue lines) for ablation study on Chain-S.
Impact of $\epsilon$.
We assess how $\epsilon$ affects the performance of $\mathsf{CARIBOU}$ and GAP, by flipping $K$ and $\epsilon$ from the previous ablation study. Specifically, we evaluate the three chain-structured datasets (Chain-S, Chain-M and Chain-L), and for each dataset, we use the same $K$ for both $\mathsf{CARIBOU}$ and GAP, and then change $\epsilon$ from $1$ to $32$.
We present the full results in Figure 9 in Appendix and one setting in Figure 5(b) ($K=10$ on Chain-S). Figure 7 in Appendix shows the accuracy of non-private $\mathsf{CARIBOU}$ and GAP(ϵ=infitalic-$\epsilon=\inf$) for comparison. When $K$ is much less than the number of nodes per chain, e.g., $K=1$ vs. 8-nodes chain, $\mathsf{CARIBOU}$ and GAP cannot realize satisfactory accuracy (both under 75%) even for non-private settings, as features from distant nodes cannot be effectively learnt. If $K$ is near to or larger than the number of nodes per chain (e.g., when $K=10$ for Chain-S, as shown in Figure 5(b)), though both GAP and $\mathsf{CARIBOU}$ see very high accuracy for non-private mode, the accuracy of GAP drops to 50% at $\epsilon=1$ and further decreases with increased $\epsilon$, suggesting the noise magnitude are not properly controlled. On the other hand, $\mathsf{CARIBOU}$ sees steady growth of accuracy along with increased $\epsilon$, which is a desired outcome for privacy protection. Notably, the accuracy of $\mathsf{CARIBOU}$ and GAP is lower than the Chain-S result in Table 3, because $K=6,7$ are chosen for best cases.
Finding 3.
Private GNNs face fundamental trade-offs between privacy, utility, and model depth $K$. Chain-structured datasets can serve as an effective benchmark for long-range scenarios due to configurable topology. Model utility becomes more susceptible to the DP noise if GNNs are tightly coupled with the underlying graph structure. Stronger privacy requires larger noise, while deeper GNN architectures amplify noise effects although privacy converges as $K\rightarrow\infty$ theoretically.
Other hyper-parameters.
Here, we evaluated the impact of other hyper-parameters in CGL, including $C_{\mathsf{L}},\alpha_{1},\beta$. For the ease of reading, we draw Figure 6 of classification accuracy using Chain-Sand Cora. As $C_{\mathsf{L}}$ constrains the features learned at each aggregation, in the relatively weak privacy guarantee ($\epsilon=16,32$), Figures 6(a) and 6(b) empirically confirms that the accuracy improves with $C_{\mathsf{L}}$ increases. In contrast, for strong privacy guarantee ($\epsilon=1,2$), larger $C_{\mathsf{L}}$ reduces the model accuracy due to the accumulated large noise. Figures 6(c) and 6(d) describe the ratio ($\alpha_{1}$) of learning from the graph, where $\alpha_{1}=1.0(\alpha_{2}=0.0)$ means the information from adjacent matrix is utilized at the maximum degree. Larger $\alpha_{1}$ leads to higher accuracy across varying $\epsilon$ in general, suggesting $\mathsf{CARIBOU}$ is able to achieve good balance between graph connectivity and privacy. The impact of $\beta$, which decides the power of residual connection between node features and CGL, is different on the two datasets. Since Chain-S is designed to tailor graph topology over node features, increasing $\beta$ to a large value (e.g., 15) might hurt accuracy. For Cora with rich node features, the model accuracy is generally increased along with $\beta$.
Figure 6. Classification accuracy under $C_{\mathsf{L}},\alpha_{1},\beta$ of CGL.
Finding 4.
All parameters $\sigma,C_{\mathsf{L}},\beta,\alpha_{1}$ in perturbed CGL contribute to the privacy-utility trade-off. Empirical analysis of $C_{\mathsf{L}},\beta,\alpha_{1}$ confirms the trade-offs by controlling the ratio of learnable node embeddings and initial node features, and topological attributes that affect the noise susceptibility.
7. Conclusion
In this study, we provide a theoretical analysis establishing a convergent privacy budget for private GNNs. Our analysis addresses a longstanding limitation in perturbed message-passing architectures — namely, the linear accumulation of noise with depth — by showing that privacy loss can remain bounded as the number of layers increases. Consequently, deeper models can be deployed with a significantly improved privacy-utility trade-off. Our analysis is broadly applicable, requiring only two conditions that are commonly satisfied in practice: the use of hidden intermediate states — a standard design choice — and contractive message passing layers, which are often observed empirically.
To demonstrate the practical implications of our theory, we introduce a novel private GNN framework, $\mathsf{CARIBOU}$, which incorporates a simple yet effective Contractive Graph Layer (CGL) that theoretically guarantees the contractiveness required by our analysis. $\mathsf{CARIBOU}$ further integrates optimized privacy budgeting, and modular auditing mechanisms to deliver strong privacy guarantees while preserving model utility. Empirical results show that $\mathsf{CARIBOU}$ substantially improves the privacy-utility trade-off and enhances robustness to membership inference attacks.
Appendix A Review on Message Passing GNNs
Every layer in the Graph Convolutional Network (GCN) (Kipf and Welling, 2017) can be expressed as a message passing layer, where the aggregation function simply computes weighted sums of the features of the neighbors. The GCN layer applied to graph ${\bm{G}}$ with a node feature matrix ${\bm{X}}^{(k)}$ can be expressed as:
(9) $\text{GCN}_{G}({\bm{X}}_{u}^{(k)})=\sigma\left(\hat{{\bm{A}}}{\bm{X}}^{(k)}{\bm{W}}^{(k)}\right),$
where $\hat{{\bm{A}}}={\bm{D}}^{-\frac{1}{2}}({\bm{A}}+{\bm{I}}){\bm{D}}^{-\frac{1}{2}}$ is the symmetrically normalized adjacency matrix of the graph, ${\bm{I}}$ is the identity matrix, ${\bm{D}}$ is the degree matrix of the graph (a diagonal matrix where $d_{ii}=\sum_{j}a_{ij}$), $a_{ij}$ is the $(i,j)$-th entry of the adjacency matrix ${\bm{A}}$, ${\bm{W}}^{(k)}$ is a learnable weight matrix at layer $k$, and $\sigma$ is a non-linear activation function.
Other simple variants include replacing the mean-type aggregation $\hat{{\bm{A}}}{\bm{X}}^{(k)}$ in GCN with a random walk adjacency matrix $\tilde{{\bm{A}}}{\bm{X}}^{(k)}$, where $\tilde{{\bm{A}}}={\bm{D}}^{-1}{\bm{A}}$ is the random walk normalized adjacency matrix. Alternative aggregation functions include sum or max aggregation, where $\hat{{\bm{A}}}{\bm{X}}^{(k)}$ is replaced with $\sum_{v\in{\mathcal{N}}(u)}{\bm{X}}_{v}^{(k)}$ (like in GAP) or $\max_{v\in{\mathcal{N}}(u)}{\bm{X}}_{v}^{(k)}$, respectively.
Appendix B Preliminary on $f$-DP and Related Definitions
In this section, we review the $f$-differential privacy (DP) framework with its definition using trade-off functions, and its special case of Gaussian differential privacy (GDP). We will also review some related results that will be used in our analysis.
The $f$-differential privacy framework (Dong et al., 2022; Kairouz et al., 2017) is based on hypothesis testing where $f$ denotes a trade-off function between type I and type II errors. Given two output distributions $P$ and $Q$ of a mechanism $\mathcal{M}$ on neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, the problem is to distinguish between them: $H_{0}:\text{dataset is }\mathcal{D}$ vs. $H_{1}:\text{dataset is }\mathcal{D}^{\prime}$. For a rejection function $\phi:\mathcal{X}\to[0,1]$, the type I error is $\mathbb{E}_{P}\phi$ and the type II error is $1-\mathbb{E}_{Q}\phi$. The privacy guarantee is formalized through the trade-off function defined below.
This function characterizes the minimum Type II error (false negative rate) as a function of the maximum Type I error (false positive rate) in hypothesis testing between $P$ and $Q$. A special case of the trade-off function is the Gaussian tradeoff function, which is used to characterize privacy in the Gaussian differential privacy (GDP) setting.
Definition 0 (Trade-off Function (Bok et al., 2024, Theorem 4.2)).
For distributions $P$, $Q$ on the same measurable space, the trade-off function $T(P,Q):[0,1]\to[0,1]$ is defined as:
[…]
A randomized algorithm $\mathcal{M}$ satisfies $f$-differential privacy if for any adjacent datasets $D$ and $D^{\prime}$, $T(\mathcal{M}(D),\mathcal{M}(D^{\prime}))\geq f$.
Note that a larger tradeoff function $f$ implies it is more difficult to distinguish the two neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, and thus the mechanism $\mathcal{M}$ is more private.
A special case of $f$-DP is the Gaussian differential privacy (GDP) (Dong et al., 2022), which is defined as follows.
Definition 0 (Gaussian Differential Privacy (Bok et al., 2024, Definition 2.1)).
A randomized algorithm $\mathcal{M}$ is $\mu$-GDP if for any adjacent datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, $T(\mathcal{M}(D),\mathcal{M}(D^{\prime}))\geq G_{\mu}$, where $G_{\mu}$ is the Gaussian tradeoff function defined as
$G_{\mu}=T(\mathcal{N}(0,1),\mathcal{N}(\mu,1)).$
More specifically, the values of $G_{\mu}$ at $\alpha\in[0,1]$ can be computed as
$G_{\mu}(\alpha)=\Phi\left(\Phi^{-1}(1-\alpha)-\mu\right),$
where $\Phi$ is the cumulative distribution function of the standard normal distribution.
A connivent tool for analyzing the composition of $f$-DP is the tensor product of trade-off functions, which is defined as follows.
Definition 0 ((Dong et al., 2022, Definition 3.1)).
The tensor product of two trade-off functions $f=T(P,Q)$ and $g=T(P^{\prime},Q^{\prime})$ is defined as
$f\otimes g=T(P\times P^{\prime},Q\times Q^{\prime}).$
The tensor product of trade-off functions satisfies the following properties as proved in Dong et al. (2022, Proposition D.1).
Lemma 0 ((Dong et al., 2022, Proposition D.1)).
The tensor product of trade-off functions satisfies the following properties:
  • $f\otimes g$ is a well-defined trade-off function.
  • $f\otimes\mathrm{Id}=f$, where $\mathrm{Id}$ is the identity function.
  • for GDP, $G_{\mu_{1}}\otimes G_{\mu_{2}}\otimes\cdots\otimes G_{\mu_{n}}=G_{\sqrt{\mu_{1}^{2}+\mu_{2}^{2}+\cdots+\mu_{n}^{2}}}$.
Lemma 0 (Post-processing inequality).
Let $P,Q$ be two probability distributions and $K$ is some map possibly random, then
(10) $T(K(P),K(Q))\geq T(P,Q).$
The following theorem shows the relationship between GDP and RDP, as well as the conversion from RDP to DP.
Theorem 6 (GDP to RDP (Bok et al., 2024, Lemma A.4)).
If a mechanism is $\mu$-GDP, then it satisfies $(\alpha,\frac{1}{2}\alpha\mu^{2})$-RDP for all $\alpha>1$.
Theorem 7 (RDP to DP (Mironov, 2017, Proposition 3)).
If $f$ is an $(\alpha,\epsilon)$-RDP mechanism, then for any $\delta\in(0,1)$, it satisfies $(\epsilon+\frac{\log(1/\delta)}{\alpha-1},\delta)$-DP.
Appendix C Proofs for results in Section 5
C.1. Proof of results in Section 5.1
Proof of Proposition 1.
[…]
Proof of Corollary 2.
The proof follows directly from applying the RDP composition Theorem 6 to Proposition 1 for $K$ independent layers. The RDP guarantee of the $K$-layer GNN is then converted to $(\epsilon,\delta)$-DP using Theorem 7. ∎
C.2. Proof of results in Section 5.2
Proof of Theorem 5.
The proof is reproduced from the original proof in Bok et al. (2024), we include it here for self-containedness and to clarify that the contractive condition (Lipschitz constant bounded by $\gamma<1$) is only required for iteration steps $k\geq 2$.
Let ${\bm{X}}^{(k)}$ and ${\bm{X}}^{\prime(k)}$ be the outputs of the two CNI processes, that is,
${\bm{X}}^{(k+1)}$ $=\Pi_{{\mathcal{K}}}\bigl{(}\phi_{k+1}({\bm{X}}^{(k)})+{\bm{Z}}^{(k+1)}\bigr{)},$
${\bm{X}}^{\prime(k+1)}$ $=\Pi_{{\mathcal{K}}}\bigl{(}\phi^{\prime}_{k+1}({\bm{X}}^{\prime(k)})+{\bm{Z}}^{\prime(k+1)}\bigr{)}.$
The key idea in the proof, following Bok et al. (2024), is to construct an auxiliary interpolating sequence $\{\widetilde{{\bm{X}}}^{(k)}\}_{k=0}^{K}$ between the two CNI processes. For each step $k$, we define:
(11) $\widetilde{{\bm{X}}}^{(k+1)}=\Pi_{{\mathcal{K}}}\bigl{(}\lambda_{k+1}\phi_{k+1}({\bm{X}}^{(k)})+(1-\lambda_{k+1})\phi^{\prime}_{k+1}(\widetilde{{\bm{X}}}^{(k)})+{\bm{Z}}^{(k+1)}\bigr{)},$
That is, $\widetilde{{\bm{X}}}^{(k+1)}$ interpolates between using $\phi_{k+1}$ and $\phi^{\prime}_{k+1}$ at each step with a mixing parameter $\lambda_{k+1}\in[0,1]$, where $\lambda_{K}=1$ so that $\widetilde{{\bm{X}}}^{(K)}={\bm{X}}^{(K)}$. Note that this interpolation process uses the same noise vector ${\bm{Z}}^{(k+1)}$ as in the original CNI process for ${\bm{X}}^{(k+1)}$.
We now recall a lemma from Bok et al. (2024) that establishes the trade-off function bound for one step of interpolation.
Lemma 0 ((Bok et al., 2024, Lemma 3.2)).
Suppose that $\phi$ and $\phi^{\prime}$ are $c$-Lipschitz functions and that $\|\phi(x)-\phi^{\prime}(x)\|\leq s$ for all $x\in{\mathcal{K}}$. Then for any $\lambda\geq 0$ and any random variable $\widetilde{{\bm{X}}}$ satisfying $\|{\bm{X}}-\widetilde{{\bm{X}}}\|\leq z$, there is
$T(\lambda\phi({\bm{X}})+(1-\lambda)\phi^{\prime}(\widetilde{{\bm{X}}})+{\mathcal{N}}(0,\sigma^{2}),\phi^{\prime}({\bm{X}}^{\prime})+{\mathcal{N}}(0,\sigma^{2}))\geq$
$\quad T(\widetilde{{\bm{X}}},{\bm{X}}^{\prime})\otimes G\left(\frac{\lambda(cz+s)}{\sigma}\right),$
With the help of this lemma, we can then bound the trade-off function between ${\bm{X}}^{(k)}$ and ${\bm{X}}^{\prime(k)}$ by first bound the distance between ${\bm{X}}^{(k)}$ and $\widetilde{{\bm{X}}}^{(k)}$ and then applying the above lemma iteratively.
We let $c_{k}$ be the maximum of the Lipschitz constants of $\phi_{k}$ and $\phi^{\prime}_{k}$ for $k=1,\dots,K$, and by our assumption, $c_{k}<1$ for $k=2,\dots,K$. We now track the distance between ${\bm{X}}^{(k)}$ and the interpolated sequence $\widetilde{{\bm{X}}}^{(k)}$.
Claim. Let $z_{k}$ be a sequence of non-negative numbers given by $z_{0}=0$ and $z_{k+1}=(1-\lambda_{k+1})(c_{k+1}z_{k}+s)$ for $k=0,\dots,K-1$. Let $\widetilde{{\bm{X}}}^{(k)}$ be the output of the interpolated CNI process in Equation 11. Then we have $\|{\bm{X}}^{(k)}-\widetilde{{\bm{X}}}^{(k)}\|\leq z_{k}$ for all $k=0,\dots,K$.
Proof of Claim. The claim is proved by induction. For $k=0$, we have $\|{\bm{X}}^{(0)}-\widetilde{{\bm{X}}}^{(0)}\|=0$ as $\widetilde{{\bm{X}}}^{(0)}={\bm{X}}^{(0)}$. For $k\geq 1$, we have
$\|{\bm{X}}^{(k)}-\widetilde{{\bm{X}}}^{(k)}\|$ $\leq\|\Pi_{{\mathcal{K}}}(\phi_{k}({\bm{X}}^{(k-1)})+{\bm{Z}}^{(k)})$
$-\Pi_{{\mathcal{K}}}(\lambda_{k}\phi_{k}({\bm{X}}^{(k-1)})+(1-\lambda_{k})\phi^{\prime}_{k}(\widetilde{{\bm{X}}}^{(k-1)})+{\bm{Z}}^{(k)})\|$
$\leq\|\phi_{k}({\bm{X}}^{(k-1)})-\lambda_{k}\phi_{k}({\bm{X}}^{(k-1)})-(1-\lambda_{k})\phi^{\prime}_{k}(\widetilde{{\bm{X}}}^{(k-1)})\|$
$\leq(1-\lambda_{k})\|\phi_{k}({\bm{X}}^{(k-1)})-\phi^{\prime}_{k}(\widetilde{{\bm{X}}}^{(k-1)})\|$
$\leq(1-\lambda_{k})(\|\phi_{k}({\bm{X}}^{(k-1)})-\phi^{\prime}_{k}(\widetilde{{\bm{X}}}^{(k-1)})\|$
$\quad+\|\phi_{k}({\bm{X}}^{(k-1)})-\phi_{k}(\widetilde{{\bm{X}}}^{(k-1)})\|)$
$\leq(1-\lambda_{k})(c_{k}z_{k-1}+s),$
where the second inequality follows from the fact that $\Pi_{{\mathcal{K}}}$ is a projection onto a convex set and hence has Lipschitz constant $1$ ((Bok et al., 2024, Lemma 2.9)). This concludes the proof of the claim by induction. We let $a_{k+1}=\lambda_{k+1}(c_{k+1}z_{k}+s)$ for $k=0,\dots,K-1$. In particular, $a_{1}=\lambda_{1}s$ regardless of the value $c_{1}$ since it is multiplied with $z_{0}=0$. This is the main our observation that the Lipchitz constant of the first iteration does not affect the constant $a_{1}$ and consequently the privacy guarantee.
We can now apply Lemma 1 iteratively to conclude the proof. There is,
$T({\bm{X}}^{(K)},{\bm{X}}^{\prime(K)})$ $=T(\widetilde{{\bm{X}}}^{(K)},{\bm{X}}^{\prime(K)})$
$\leq T(\lambda_{K}\phi_{K}({\bm{X}}^{(K-1)})+(1-\lambda_{K})\phi^{\prime}_{K}(\widetilde{{\bm{X}}}^{(K-1)})+{\bm{Z}}^{(K)},$
$\quad\phi^{\prime}_{K}({\bm{X}}^{\prime(K-1)})+{\bm{Z}}^{(K)})$
$\leq T(\widetilde{{\bm{X}}}^{(K-1)},{\bm{X}}^{\prime(K-1)})\otimes G\left(\frac{a_{K}}{\sigma}\right)$
$\leq T(\widetilde{{\bm{X}}}^{(K-2)},{\bm{X}}^{\prime(K-1)})\otimes G\left(\frac{a_{K-1}}{\sigma}\right)\otimes G\left(\frac{a_{K}}{\sigma}\right)$
$\leq T(\widetilde{{\bm{X}}}^{(0)},{\bm{X}}^{\prime(0)})\otimes G\left(\frac{a_{1}}{\sigma}\right)\otimes G\left(\frac{a_{2}}{\sigma}\right)\cdots G\left(\frac{a_{K}}{\sigma}\right).$
$\leq G\left(\frac{1}{\sigma}\sqrt{\sum_{k=1}^{K}a_{k}^{2}}\right)$
where the last inequality follows from the fact that $\widetilde{{\bm{X}}}^{(0)}={\bm{X}}^{(0)}$ and property of Gaussian tradeoff function in Lemma 4.
Lastly, by the assumption that $\phi_{k}$ and $\phi^{\prime}_{k}$ are $\gamma$-Lipschitz for $k=2,\dots,K$, we can let $c_{k}=\gamma$ for $k=2,\dots,K$. We set $\lambda_{k}=\frac{\gamma^{K-k}(1-\gamma^{2})}{1-\gamma^{K-k+2}-\gamma^{k}+\gamma^{K}}$ for $k=1,\dots,K$ as computed in Bok et al. (2024, Lemma C.6). Then $z_{k}=\frac{(1-\gamma^{k})(1-\gamma^{K-k})}{(1+\gamma^{K})(1-\gamma)}s$, and consequently, $a_{k}=\frac{\gamma^{K-k}(1+\gamma)}{1+\gamma^{k}}s$. In this case, $\sum_{k=1}^{K}a_{k}^{2}=\frac{(1-\gamma^{K})(1+\gamma)}{(1+\gamma^{K})(1-\gamma)}s^{2}$. This concludes the proof of the theorem. ∎
Proof of Theorem 6.
Let $G,G^{\prime}$ be two adjacent graphs dataset. Let $\mathrm{MP}_{G}$ represent message passing operation for a fixed graph $G$, with projection $\Pi_{\mathcal{K}}$ ensuring output lies in convex set ${\mathcal{K}}$. Then starting from ${\bm{X}}^{(0)}$, the perturbed message passing GNN with $\mathrm{MP}$ can be represented as the CNI process
$\text{CNI}({\bm{X}}^{(0)},\{\mathrm{MP}_{G}\}_{k=1}^{K},\{\mathcal{N}(0,\sigma^{2}{\bm{I}}_{d})\}_{k=1}^{K},{\mathcal{K}}).$
Similarly, the perturbed message passing GNN with $\mathrm{MP}^{\prime}$ for graph $G^{\prime}$ is represented as
$\text{CNI}({\bm{X}}^{(0)},\{\mathrm{MP}_{{\bm{G}}^{\prime}}\}_{k=1}^{K},\{\mathcal{N}(0,\sigma^{2}{\bm{I}}_{d})\}_{k=1}^{K},{\mathcal{K}}).$
By assumptions, both $\mathrm{MP}_{G}$ and $\mathrm{MP}_{{\bm{G}}^{\prime}}$ are contractive with Lipschitz constant $\gamma<1$ for layers $k=2,\dots,K$ and sensitivity bound of $\mathrm{MP}$ implies that $\|\mathrm{MP}_{G}(x)-\mathrm{MP}_{{\bm{G}}^{\prime}}(x)\|\leq\Delta(\mathrm{MP})$ for all $x\in{\mathcal{K}}$. This shows that we can the meta-theorem in Theorem 5 to bound the trade-off function between the two CNI processes as
$T({\bm{X}}^{(K)},{\bm{X}}^{\prime(K)})$ $\geq G\left(\frac{\Delta(\mathrm{MP})}{\sigma}\sqrt{\frac{1-\gamma^{K}}{1+\gamma^{K}}\frac{1+\gamma}{1-\gamma}}\right)$
This implies that the $K$-layer perturbed message passing GNN with $\mathrm{MP}$ is $\frac{\Delta(\mathrm{MP})}{\sigma}\sqrt{\frac{1-\gamma^{K}}{1+\gamma^{K}}\frac{1+\gamma}{1-\gamma}}$-Gaussian differential privacy (GDP) defined in Definition 2. By applying Lemma 6, we then obtain the stated RDP guarantee. Then by applying Theorem 7, we can convert the RDP guarantee to $(\epsilon,\delta)$-DP. ∎
C.3. Proof of results in Section 5.3
Proof of Proposition 7.
Let ${\bm{Y}}^{(k-1)},{\bm{Y}}^{\prime(k-1)}\in{\mathcal{K}}$ be two inputs to the message passing operator $\mathrm{MP}_{G}$ at layer $k$. Since $k\geq 2$, the residue term $\beta{\bm{X}}^{(0)}$ is independent of the input ${\bm{Y}}^{(k-1)}$ and ${\bm{Y}}^{\prime(k-1)}$, and thus does not affect the Lipschitz constant. We can write the difference between the outputs of $\mathsf{CGL}$ as follows:
$\|\mathsf{CGL}({\bm{Y}}^{(k-1)})-\mathsf{CGL}({\bm{Y}}^{\prime(k-1)})\|$
$\leq\|C_{\mathsf{L}}(\alpha_{1}\hat{{\bm{A}}}{\bm{Y}}^{(k-1)}+\alpha_{2}\mathsf{Mean}({\bm{Y}}^{(k-1)}))+\beta{\bm{X}}^{(0)}$
$\quad\quad-C_{\mathsf{L}}(\alpha_{1}\hat{{\bm{A}}}{\bm{Y}}^{\prime(k-1)}+\alpha_{2}\mathsf{Mean}({\bm{Y}}^{\prime(k-1)}))-\beta{\bm{X}}^{(0)}\|$
$\leq C_{\mathsf{L}}\|\alpha_{1}(\hat{{\bm{A}}}{\bm{Y}}^{(k-1)}-\hat{{\bm{A}}}{\bm{Y}}^{\prime(k-1)})+\alpha_{2}(\mathsf{Mean}({\bm{Y}}^{(k-1)})-\mathsf{Mean}({\bm{Y}}^{\prime(k-1)}))\|$
$\leq C_{\mathsf{L}}(\alpha_{1}+\alpha_{2})\|{\bm{Y}}^{(k-1)}-{\bm{Y}}^{\prime(k-1)}\|=C_{\mathsf{L}}\|{\bm{Y}}^{(k-1)}-{\bm{Y}}^{\prime(k-1)}\|,$
where the second line follows from the fact that the operator norms of $\hat{{\bm{A}}}$ and $\mathsf{Mean}$ are bounded by $1$. ∎
Proof of Theorem 8.
Let ${\bm{G}},{\bm{G}}^{\prime}$ be two edge adjacent graphs and $\hat{{\bm{A}}},\hat{{\bm{A}}}^{\prime}$ be the corresponding adjacency matrices of ${\bm{G}},{\bm{G}}^{\prime}$ respectively. Without loss of generality, we assume that the edge $e_{uv}$ is added to ${\bm{G}}$ to form ${\bm{G}}^{\prime}$ for two nodes $u$ and $v$. Then the $\mathsf{CGL}$ layer updates the node features as follows:
${\bm{X}}^{(k)}$ $=C_{\mathsf{L}}(\alpha_{1}\hat{{\bm{A}}}{\bm{X}}^{(k-1)}+\alpha_{2}\mathsf{Mean}({\bm{X}}^{(k-1)}))+\beta{\bm{X}}^{(0)},$
${\bm{X}}^{\prime(k)}$ $=C_{\mathsf{L}}(\alpha_{1}\hat{{\bm{A}}}^{\prime}{\bm{X}}^{\prime(k-1)}+\alpha_{2}\mathsf{Mean}({\bm{X}}^{\prime(k-1)}))+\beta{\bm{X}}^{(0)}.$
The difference between the two outputs is given by the aggregation of $\hat{{\bm{A}}}$ and $\hat{{\bm{A}}}^{\prime}$. Then the edge-level sensitivity $\Delta_{e}(\mathsf{CGL})$ is amount to bound $\|\hat{{\bm{A}}}{\bm{X}}^{(k)}-\hat{{\bm{A}}}^{\prime}{\bm{X}}^{\prime(k)}\|_{F}$. Since only one edge is added, the difference between $\hat{{\bm{A}}}$ and $\hat{{\bm{A}}}^{\prime}$ is only on the row corresponding to $u$ and $v$.
For row $u$, we need to bound $\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$. For $(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}$, we can write it as
(12) $(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}$ $=\frac{1}{d_{u}+1}{\bm{X}}^{(k)}_{u}+\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+1}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}$
where $d_{u}$ is the degree of node $u$ in graph ${\bm{G}}$ and ${\bm{N}}_{u}$ is the neighbors of node $u$ in graph ${\bm{G}}$. For $(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}$, with the same notation for $d_{u}$ and ${\bm{N}}_{u}$, we can write it as
(13) $(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}$ $=\frac{1}{d_{u}+2}{\bm{X}}^{(k)}_{u}+\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}$
$+\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{v}^{\prime}+1}}{\bm{X}}^{(k)}_{v}$
where $d^{\prime}_{v}$ is the degree of node $v$ in graph ${\bm{G}}^{\prime}$.
Then there is
$\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$
$\leq\left\|\frac{1}{d_{u}+1}{\bm{X}}^{(k)}_{u}-\frac{1}{d_{u}+2}{\bm{X}}^{(k)}_{u}\right\|_{2}$
$+\quad\left\|\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+1}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}-\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}\right\|_{2}$
$+\left\|\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{v}^{\prime}+1}}{\bm{X}}^{(k)}_{v}\right\|_{2}$
$\leq\frac{1}{(d_{u}+1)(d_{u}+2)}+\sum_{w\in N_{u}}\frac{1}{\sqrt{d_{w}+1}}(\frac{1}{\sqrt{d_{u}+1}}-\frac{1}{\sqrt{d_{u}+2}})$
$+\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{v}^{\prime}+1}}$
$\leq\frac{1}{(d_{u}+1)(d_{u}+2)}+\frac{d_{u}}{\sqrt{d_{w}+1}}(\frac{1}{\sqrt{d_{u}+1}}-\frac{1}{\sqrt{d_{u}+2}})$
$+\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{v}^{\prime}+1}}$
To bound $\frac{d_{u}}{\sqrt{d_{w}+1}}(\frac{1}{\sqrt{d_{u}+1}}-\frac{1}{\sqrt{d_{u}+2}})$, we study the monotonicity of the function $f(x)=\frac{x}{\sqrt{x+1}}-\frac{x}{\sqrt{x+2}}$ for $x>0$. It turns out that $f(x)$ only has one positive critical point and is around $x=2.9$, and when evaluated on integers, $f(x)$ increases from $x=1$ to $x=3$ and then decreases from $x=3$ to $\infty$. Thus, when the minimum degree $D_{\min}$ of ${\bm{G}}$ is larger than $3$, the function $f(x)$ is maximized at $x=D_{\min}$, and we have
$\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$
$\leq\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{D_{\min}}{\sqrt{D_{\min}+1}}(\frac{1}{\sqrt{D_{\min}+1}}-\frac{1}{\sqrt{D_{\min}+2}})$
$+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}$
where we use the fact that the minimum degree of ${\bm{G}}^{\prime}$ is larger than that of ${\bm{G}}$. When $1\leq D_{\min}\leq 3$, we can bound the function $f(x)$ by $f(3)=\frac{3}{\sqrt{4}}-\frac{3}{\sqrt{5}}$. This results in
$\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$
$\leq\frac{1}{(D_{\min}+1)(D_{\min}+2)}+(\frac{3}{\sqrt{4}}-\frac{3}{\sqrt{5}})\frac{1}{\sqrt{D_{\min}+1}}$
$\quad\quad+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}$
For notation convenience, we use $C(D_{\min})$ to denote the piecewise function of $D_{\min}$, which is defined as
(14) $C(D_{\min})=\begin{cases}\frac{D_{\min}}{\sqrt{D_{\min}+1}}-\frac{D_{\min}}{\sqrt{D_{\min}+2}}&D_{\min}>3\\ (\frac{3}{\sqrt{4}}-\frac{3}{\sqrt{5}})&1\leq D_{\min}\leq 3\\ \end{cases}$
Therefore, the affect of modifying an edge on a single node $u$ of $\hat{{\bm{A}}}{\bm{X}}^{(k)}$ is bounded by
(15) $\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$
$\leq\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}$
The same analysis can be applied to the row $v$ of $\hat{{\bm{A}}}{\bm{X}}^{(k)}$ and $\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)}$. The edge sensitivity $\Delta_{e}(\mathsf{CARIBOU})$ can then be bounded as the following:
$\Delta_{e}(\mathsf{CARIBOU}):=\max_{{\bm{G}},{\bm{G}}^{\prime}}\|\hat{{\bm{A}}}{\bm{X}}^{(k)}-\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)}\|_{F}$
$\leq\alpha_{1}C_{\mathsf{L}}\sqrt{\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}^{2}+\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{v}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{v}\|_{2}^{2}}$
$\leq\sqrt{2}\alpha_{1}C_{\mathsf{L}}\left(\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{D_{\min}+1}}\right)$
Proof of Theorem 9.
Let ${\bm{G}},{\bm{G}}^{\prime}$ be two node adjacent graphs and $\hat{{\bm{A}}},\hat{{\bm{A}}}^{\prime}$ be the corresponding adjacency matrices of ${\bm{G}},{\bm{G}}^{\prime}$ respectively. Without loss of generality, we assume that the node $v$ is added to ${\bm{G}}$ to form ${\bm{G}}^{\prime}$ and connected to nodes $N_{v}$ in ${\bm{G}}$. The layer updates the node features as follows:
${\bm{X}}^{(k)}$ $=C_{\mathsf{L}}(\alpha_{1}\hat{{\bm{A}}}{\bm{X}}^{(k-1)}+\alpha_{2}\mathsf{Mean}({\bm{X}}^{(k-1)}))+\beta{\bm{X}}^{(0)},$
${\bm{X}}^{\prime(k)}$ $=C_{\mathsf{L}}(\alpha_{1}\hat{{\bm{A}}}^{\prime}{\bm{X}}^{\prime(k-1)}+\alpha_{2}\mathsf{Mean^{\prime}}({\bm{X}}^{\prime(k-1)}))+\beta{\bm{X}}^{(0)}.$
The difference between the two outputs is given by the aggregation of $\hat{{\bm{A}}}$ and $\hat{{\bm{A}}}^{\prime}$ as well as the mean operator $\mathsf{Mean}$ and $\mathsf{Mean^{\prime}}$ since ${\bm{G}}^{\prime}$ has one more node than ${\bm{G}}$. Then the node-level sensitivity $\Delta_{v}(\mathsf{CGL})$ can be bounded as follows:
$\Delta_{n}(\mathsf{CGL})$
$=\max_{{\bm{G}},{\bm{G}}^{\prime}}\|\mathsf{CGL}({\bm{X}}^{(k)})-\mathsf{CGL}^{\prime}({\bm{X}}^{(k)})\|_{F}$
$\leq\|{\bm{X}}_{v}^{\prime}\|_{2}+\sum_{u\in{\bm{N}}_{v}}\alpha_{1}C_{\mathsf{L}}\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$
$+\quad\sum_{u\in{\bm{N}}_{v}}\alpha_{2}C_{\mathsf{L}}\|\mathsf{Mean}({\bm{X}}^{(k)})_{u}-\mathsf{Mean^{\prime}}({\bm{X}}^{(k)})_{u}\|_{2}$
For the first term $\|{\bm{X}}_{v}^{\prime}\|_{2}$, it is bounded by $1$ by constraint of ${\mathcal{K}}$. For the second term, we can argue similar as in the proof of Theorem 8. For nodes $u\in{\bm{N}}_{v}$, there is
$\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$
$\leq\left\|\frac{1}{d_{u}+1}{\bm{X}}^{(k)}_{u}-\frac{1}{d_{u}+2}{\bm{X}}^{(k)}_{u}\right\|_{2}$
$+\quad\left\|\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+1}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}-\sum_{w\in{\bm{N}}_{u}}\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{w}+1}}{\bm{X}}^{(k)}_{w}\right\|_{2}$
$+\left\|\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{v}^{\prime}+1}}{\bm{X}}^{(k)}_{v}\right\|_{2}$
$\leq\frac{1}{(d_{u}+1)(d_{u}+2)}+\sum_{w\in N_{u}}\frac{1}{\sqrt{d_{w}+1}}(\frac{1}{\sqrt{d_{u}+1}}-\frac{1}{\sqrt{d_{u}+2}})$
$+\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{v}^{\prime}+1}}$
$\leq\frac{1}{(d_{u}+1)(d_{u}+2)}+\frac{d_{u}}{\sqrt{d_{w}+1}}(\frac{1}{\sqrt{d_{u}+1}}-\frac{1}{\sqrt{d_{u}+2}})$
$+\frac{1}{\sqrt{d_{u}+2}\sqrt{d_{v}^{\prime}+1}}$
$\leq\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}$
Then the summation term $\sum_{u\in{\bm{N}}_{v}}\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$ can be bounded by
$\sum_{u\in{\bm{N}}_{v}}\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$
$\leq\sum_{u\in{\bm{N}}_{v}}\left(\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}\right)$
$\leq|d_{v}^{\prime}|\left(\frac{1}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}\right)$
$\leq\frac{\sqrt{d_{v}^{\prime}}}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})\sqrt{d_{v}^{\prime}}}{\sqrt{D_{\min}+1}}+\frac{\sqrt{d_{v}^{\prime}}}{\sqrt{D_{\min}+2}\sqrt{d_{v}^{\prime}+1}}$
$\leq\frac{\sqrt{d_{v}^{\prime}}}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})\sqrt{d_{v}^{\prime}}}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}}$
$\leq\frac{\sqrt{D_{\max}}}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})\sqrt{D_{\max}}}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}},$
where $D_{\max}$ is the maximum degree of the graph ${\bm{G}}$ and $d_{v}^{\prime}$ is the degree of node $v$ in graph ${\bm{G}}^{\prime}$. Additionally, there is
$\|\mathsf{Mean}({\bm{X}}^{(k)})_{u}-\mathsf{Mean^{\prime}}({\bm{X}}^{(k)})_{u}\|_{2}$
$=\left\|\frac{1}{|V|}\sum_{w\in V}{\bm{X}}^{(k)}_{w}-\frac{1}{|V|+1}\sum_{w\in V}{\bm{X}}^{(k)}_{w}-\frac{1}{|V|+1}{\bm{X}}^{(k)}_{v^{\prime}}\right\|_{2}$
$\leq\left\|\frac{1}{|V|(|V|+1)}\sum_{w\in V}{\bm{X}}^{(k)}_{w}-\frac{1}{|V|+1}{\bm{X}}^{(k)}_{v^{\prime}}\right\|_{2}$
$\leq\frac{2}{|V|+1},$
where $|V|$ is the number of nodes in graph ${\bm{G}}$.
Then for the node-level sensitivity of one layer of $\mathsf{CGL}$, we have
$\Delta_{n}(\mathsf{CGL})$
$=\max_{{\bm{G}},{\bm{G}}^{\prime}}\|\mathsf{CGL}({\bm{X}}^{(k)})-\mathsf{CGL}^{\prime}({\bm{X}}^{(k)})\|_{F}$
$\leq\|{\bm{X}}_{v}^{\prime}\|_{2}+\sum_{u\in{\bm{N}}_{v}}\alpha_{1}C_{\mathsf{L}}\|(\hat{{\bm{A}}}{\bm{X}}^{(k)})_{u}-(\hat{{\bm{A}}}^{\prime}{\bm{X}}^{(k)})_{u}\|_{2}$
$+\quad\sum_{u\in{\bm{N}}_{v}}\alpha_{2}C_{\mathsf{L}}\|\mathsf{Mean}({\bm{X}}^{(k)})_{u}-\mathsf{Mean^{\prime}}({\bm{X}}^{(k)})_{u}\|_{2}$
$\leq 1+\alpha_{1}C_{\mathsf{L}}\left(\frac{\sqrt{D_{\max}}}{(D_{\min}+1)(D_{\min}+2)}+\frac{C(D_{\min})\sqrt{D_{\max}}}{\sqrt{D_{\min}+1}}+\frac{1}{\sqrt{D_{\min}+2}}\right)$
$+\quad\alpha_{2}C_{\mathsf{L}}\frac{2|V|}{|V|+1}$
Proof of Corollary 10.
The proof follows from plugging the edge-level and node-level sensitivity of $\mathsf{CARIBOU}$ into Theorem 6. ∎
Proof of Theorem 11.
The result comes from directly applying the composition theorem of DP to different modules of $\mathsf{CARIBOU}$. ∎
Appendix D Experimental Setup
This section presents experiments setups and counterpart baselines in detail.
D.1. Datasets
Table 5 lists statistics of these datasets, including Amazon co-purchase networks (Photo, Computers (Shchur et al., 2018)), social network (Facebook (Traud et al., 2011)), and citation networks (Cora, PubMed (Sen et al., 2008)). These datasets are also widely adopted as benchmark datasets to evaluate various GNNs (Gu et al., 2020; Sajadmanesh et al., 2023; Chien et al., 2023; Yao et al., 2023). In addition, we use the synthetic graph generation algorithm developed in IGNN (Gu et al., 2020), to produce chain-structured datasets for convenient and generalized verification of multi-hopo aggregations. The chain-structured dataset can be configured with various number of nodes per chain, number of chains per class, and the number of class. In Table 6, the statistics of the four synthesized chain-structured datasets, namely Chain-S, Chain-M, Chain-L, and Chain-X  are shown. We unify the feature dimension to $5$ and focus on the structure of chain-structured dataset, and we configure the number of nodes to $\{8,10,15\}$ per chain, and the number of chains to $\{3,5\}$ per class.
Table 5. Standard Graph Dataset Statistics. “#Tra” and “#Test” are the ratios of the total nodes used for training and testing.
Dataset Node Edge Feature Class #Tra #Test
Computers $13,471$ $491,722$ $767$ $10$ $10\%$ $20\%$
Facebook $26,406$ $2,117,924$ $501$ $6$ $10\%$ $20\%$
PubMed $19,717$ $88,648$ $500$ $3$ $10\%$ $20\%$
Cora $2,708$ $10,556$ $1,433$ $7$ $10\%$ $20\%$
Photo $7,535$ $238,162$ $745$ $8$ $10\%$ $20\%$
Table 6. Chain-structured Datasets Statistics. Each chain has the same number of nodes and “Node” is the sum of nodes across chains.
Dataset Node Chain Feature Class #Tra #Test
Chain-S $48$ $6$ $5$ $2$ $8$ $32$
Chain-M $60$ $6$ $5$ $2$ $10$ $40$
Chain-L $90$ $6$ $5$ $2$ $15$ $60$
Chain-X $150$ $10$ $5$ $2$ $25$ $100$
D.2. Baseline Methods
Edge-level private algorithms.
The evaluated edge-level private algorithms are introduced below.
  • DPDGC: This algorithm decouples the neighborhood aggregation process from the transformation of node features, thereby avoiding direct aggregation of sensitive data. This decoupling enables improved privacy guarantees through the DP composition theorem, allowing for a more efficient balance between privacy guarantee and model performance.
  • GAP: This algorithm preserves edge privacy via aggregation perturbation, i.e., adding calibrated Gaussian noise to the output of the aggregation function for hiding the presence of a particular edge. GAP’s architecture involves pre-training encoder, aggregation module, and classification module, so that GAP  can reduce the privacy costs of the perturbed aggregations by one-time computation over lower-dimensional embeddings.
  • SAGE: This algorithm adopts the graph perturbation, building on the popular GraphSAGE architecture as its backbone GNN model. To realize graph perturbation, SAGE perturbs the adjacency matrix of graph using the asymmetric randomized response.
  • MLP: Typical MLP model is trained over node features, without referring to graph edges. Thus, $\epsilon=0$ always holds for MLP, revealing nothing about edges and providing complete edge-level privacy.
Vanilla algorithms.
The non-private versions of all private algorithms above are used to quantify the accuracy loss of corresponding EDP algorithms.
Appendix E More Experimental Results
E.1. Computational Overhead
Table 7. Execution Time of $3$ Training-Test Runs (seconds)
Dataset $\mathsf{CARIBOU}$ DPDGC GAP SAGE MLP Chain-S $2.66$ $3.70$ $5.15$ $2.97$ $2.34$ Chain-M $2.66$ $3.70$ $5.24$ $2.98$ $2.34$ Chain-L $2.65$ $3.69$ $5.31$ $3.00$ $2.35$ Chain-X $2.67$ $3.73$ $5.27$ $2.97$ $2.36$ Computers $3.34$ $4.47$ $6.19$ $3.41$ $2.56$ Facebook $4.01$ $6.04$ $6.30$ $3.74$ $2.74$ PubMed $3.10$ $4.63$ $6.32$ $3.34$ $2.52$ Cora $2.79$ $4.02$ $5.52$ $3.04$ $2.40$ Photo $3.04$ $3.97$ $5.93$ $3.18$ $2.48$
We measure the latency and memory usage to assess the overhead of $\mathsf{CARIBOU}$. For latency, we use the total execution time (training and testing), and in Table 7 we list the average of running $3$ times. We found $\mathsf{CARIBOU}$ has less latency than the other method (2.99 seconds of $\mathsf{CARIBOU}$ averaged among datasets, comparing to 4.22 seconds of DPDGC, 5.69 seconds of GAP, 3.18 seconds of SAGE), except MLP (2.45 seconds), which does not incur any graph operations. Notable latency increase is observed on the Facebook dataset because its large number of nodes (26,406) and edges (2,117,924), and $\mathsf{CARIBOU}$ shows a larger lead comparing the other private GNNs: 4.01, 6.04 and 6.30 for $\mathsf{CARIBOU}$, GAP and p̣ d g c.
For memory, we measure the max memory usage over the 3 runs and report the results in Table 8. For chain-structured datasets, the memory consumption averaged across datasets of $\mathsf{CARIBOU}$ (1.09MB) is close to SAGE (1.12MB) and MLP (1.05MB), while much lower than GAP (1.55MB) and DPDGC (1.55MB). On the other hand, we found higher memory consumption of $\mathsf{CARIBOU}$ is introduced on the other datasets, comparing to the baseline methods due to graph loading mechanism in GAP. Still, the memory consumption of $\mathsf{CARIBOU}$ is reasonable, which is smaller than our GPU memory limit (24GB) for all datasets. We argue the privacy, and utility benefits of $\mathsf{CARIBOU}$ outweigh its memory costs here.
Table 8. Max Memory Usage (MB)
Dataset $\mathsf{CARIBOU}$ DPDGC GAP SAGE MLP Chain-S $1.08$ $1.46$ $1.39$ $1.09$ $1.05$ Chain-M $1.08$ $1.49$ $1.44$ $1.10$ $1.05$ Chain-L $1.09$ $1.56$ $1.56$ $1.12$ $1.06$ Chain-X $1.10$ $1.68$ $1.80$ $1.16$ $1.07$ Computers $314.04$ $197.21$ $138.39$ $116.85$ $134.13$ Facebook $525.18$ $494.07$ $315.82$ $268.41$ $275.01$ PubMed $272.70$ $202.67$ $190.10$ $97.45$ $111.17$ Cora $104.94$ $56.14$ $27.53$ $32.99$ $42.19$ Photo $170.28$ $110.60$ $77.54$ $62.69$ $72.29$
E.2. Accuracy of $\mathsf{CARIBOU}$-NDP
Table 10 presents the top accuracy for NDP implemented by $\mathsf{CARIBOU}$ on 6 datasets. The results demonstrate that accuracy increases in general (except Chain-S at $\epsilon=16,32$) with larger $\epsilon$ values, aligning with the EDP results shown in Table 3. As established in Theorem 9, NDP requires injecting more noise compared to EDP under the same privacy budget, hence, the accuracy of NDP is often lower than EDP.
Regarding the other baselines, we we did not implement the NDP version of GAP because it does not correctly account for node-level sensitivity, as pointed out by Xiang et al. (Xiang et al., 2024). Other baselines like DPDGC do not aim at multi-hop GNN training, i.e., with a fixed number of layers, thus, we do not test the NDP versions for them.
Table 9. Mean accuracy over 3 runs for EDP. The best accuracy and the second-best accuracy are highlighted, respectively. The symbol
[…]
Dataset Chain-S Chain-M Chain-L Chain-X Computers Facebook PubMed Cora Photo $\epsilon=1$ $\mathsf{CARIBOU}$ $78.1\%$ $80.0\%$ $70.0\%\mathrel{\hskip 5.12128pt}$ $65.0\%\mathrel{\hskip 5.12128pt}$ $92.0\%$ $74.0\%\mathrel{\hskip 5.12128pt}$ $87.9\%$ $84.3\%$ $95.6\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.0\%$ $60.6\%\mathrel{\hskip 5.12128pt}$ $88.3\%$ $75.5\%$ $92.5\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $67.5\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $55.5\%\mathrel{\hskip 5.12128pt}$ $87.0\%$ $68.3\%\mathrel{\hskip 5.12128pt}$ $87.2\%$ $76.0\%$ $92.8\%$ SAGE $53.1\%$blacktriangledown $45.0\%$blacktriangledown $53.3\%$blacktriangledown $51.0\%$blacktriangledown $77.8\%$ $48.2\%$blacktriangledown $85.0\%$ $60.0\%$ $82.4\%$ $\epsilon=2$ $\mathsf{CARIBOU}$ $90.6\%$ $82.5\%$ $71.7\%$ $68.0\%$ $92.0\%$ $73.8\%\mathrel{\hskip 5.12128pt}$ $89.1\%$ $86.0\%$ $95.7\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.2\%$ $66.7\%\mathrel{\hskip 5.12128pt}$ $88.2\%$ $77.5\%$ $93.3\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $67.5\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $55.0\%$blacktriangledown $88.0\%$ $71.7\%\mathrel{\hskip 5.12128pt}$ $87.3\%$ $76.8\%$ $93.3\%$ SAGE $43.8\%$blacktriangledown $45.0\%$blacktriangledown $53.3\%$blacktriangledown $51.0\%$blacktriangledown $76.1\%$ $48.1\%$blacktriangledown $84.6\%$ $60.1\%$ $82.4\%$ $\epsilon=4$ $\mathsf{CARIBOU}$ $90.6\%$ $82.5\%$ $71.7\%$ $65.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $73.9\%\mathrel{\hskip 5.12128pt}$ $89.5\%$ $86.7\%$ $95.8\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.9\%$ $73.3\%\mathrel{\hskip 5.12128pt}$ $88.4\%$ $75.1\%$ $94.2\%$ GAP $62.5\%\mathrel{\hskip 5.12128pt}$ $60.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $55.5\%\mathrel{\hskip 5.12128pt}$ $88.8\%$ $73.6\%\mathrel{\hskip 5.12128pt}$ $87.7\%$ $76.9\%$ $93.8\%$ SAGE $50.0\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $54.0\%$blacktriangledown $79.1\%$ $50.3\%$blacktriangledown $85.8\%$ $63.3\%$ $85.7\%$ $\epsilon=8$ $\mathsf{CARIBOU}$ $81.2\%$ $82.5\%$ $73.3\%$ $68.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $74.2\%\mathrel{\hskip 5.12128pt}$ $89.7\%$ $87.6\%$ $95.8\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $89.4\%$ $78.6\%\mathrel{\hskip 5.12128pt}$ $88.6\%$ $76.0\%$ $94.6\%$ GAP $59.4\%\mathrel{\hskip 5.12128pt}$ $60.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $55.5\%\mathrel{\hskip 5.12128pt}$ $89.6\%$ $75.0\%\mathrel{\hskip 5.12128pt}$ $88.0\%$ $78.2\%$ $94.6\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $61.0\%\mathrel{\hskip 5.12128pt}$ $87.9\%$ $75.6\%\mathrel{\hskip 5.12128pt}$ $84.8\%$ $75.7\%$ $92.2\%$ $\epsilon=16$ $\mathsf{CARIBOU}$ $87.5\%$ $85.0\%$ $70.0\%\mathrel{\hskip 5.12128pt}$ $68.0\%$ $92.1\%$ $74.3\%\mathrel{\hskip 5.12128pt}$ $89.8\%$ $88.0\%$ $95.9\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $40.0\%$blacktriangledown $90.4\%$ $81.2\%\mathrel{\hskip 5.12128pt}$ $88.9\%$ $77.7\%$ $93.8\%$ GAP $59.4\%\mathrel{\hskip 5.12128pt}$ $70.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $54.3\%$blacktriangledown $90.0\%$ $76.0\%\mathrel{\hskip 5.12128pt}$ $88.5\%$ $80.3\%$ $94.6\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $51.0\%$blacktriangledown $90.9\%$ $79.5\%\mathrel{\hskip 5.12128pt}$ $87.6\%$ $84.7\%$ $94.1\%$ $\epsilon=32$ $\mathsf{CARIBOU}$ $90.6\%$ $82.5\%$ $78.3\%$ $67.0\%$ $92.2\%$ $73.9\%\mathrel{\hskip 5.12128pt}$ $89.8\%$ $88.2\%$ $95.8\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $40.0\%$blacktriangledown $91.3\%$ $82.9\%\mathrel{\hskip 5.12128pt}$ $88.8\%$ $79.9\%$ $94.3\%$ GAP $59.4\%\mathrel{\hskip 5.12128pt}$ $65.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $54.3\%$blacktriangledown $90.2\%$ $76.5\%\mathrel{\hskip 5.12128pt}$ $88.7\%$ $82.2\%$ $94.6\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $51.0\%$blacktriangledown $90.6\%$ $79.9\%\mathrel{\hskip 5.12128pt}$ $86.9\%$ $85.2\%$ $94.4\%$ Plain $\mathsf{CARIBOU}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $78.1\%\mathrel{\hskip 5.12128pt}$ $89.9\%$ $88.4\%$ $95.8\%$ DPDGC $59.4\%\mathrel{\hskip 5.12128pt}$ $77.5\%\mathrel{\hskip 5.12128pt}$ $63.3\%\mathrel{\hskip 5.12128pt}$ $73.0\%\mathrel{\hskip 5.12128pt}$ $92.8\%$ $86.4\%\mathrel{\hskip 5.12128pt}$ $88.1\%$ $83.9\%$ $96.2\%$ GAP $97.9\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $93.3\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $90.9\%$ $78.3\%\mathrel{\hskip 5.12128pt}$ $89.1\%$ $85.1\%$ $95.2\%$ SAGE $59.4\%\mathrel{\hskip 5.12128pt}$ $55.0\%$blacktriangledown $60.0\%\mathrel{\hskip 5.12128pt}$ $58.0\%\mathrel{\hskip 5.12128pt}$ $91.6\%$ $79.7\%\mathrel{\hskip 5.12128pt}$ $87.0\%$ $82.1\%$ $94.2\%$ MLP $40.6\%$blacktriangledown $47.5\%$blacktriangledown $46.7\%$blacktriangledown $51.0\%$blacktriangledown $83.9\%$ $51.1\%$blacktriangledown $84.9\%$ $72.9\%$ $91.1\%$
Table 10. Top Accuracy of Node-level Privacy.
Dataset Chain-S Chain-M Chain-L Chain-X Cora Photo
$\epsilon=1$ $75.0\%$ $70.0\%$ $60.0\%$ $60.0\%$ $66.8\%$ $93.7\%$
$\epsilon=2$ $78.1\%$ $70.0\%$ $61.7\%$ $61.0\%$ $70.3\%$ $94.7\%$
$\epsilon=4$ $78.1\%$ $70.0\%$ $63.3\%$ $61.0\%$ $75.5\%$ $95.0\%$
$\epsilon=8$ $78.1\%$ $72.5\%$ $63.3\%$ $62.0\%$ $79.9\%$ $95.3\%$
$\epsilon=16$ $75.0\%$ $72.5\%$ $68.3\%$ $63.0\%$ $82.1\%$ $95.3\%$
$\epsilon=32$ $71.9\%$ $72.5\%$ $73.3\%$ $64.0\%$ $84.7\%$ $95.5\%$
E.3. More Results of EDP
Table 9 shows the mean accuracy compared between $\mathsf{CARIBOU}$ and other baselines for all datasets.
Figure 7 demonstrates the learning behavior on chain-structured datasets, where increasing the number of aggregation hops ($K$) leads to continuous improvement in accuracy for both GAP and $\mathsf{CARIBOU}$ in general. This indicates that deeper message passing is beneficial for capturing information on long chains. For the first $15$ hops, $\mathsf{CARIBOU}$ outperforms GAP, suggesting it is more effective to aggregate node features at smaller range.
Figure 7. Accuracy of non-private $\mathsf{CARIBOU}$ and GAP under different $K$.
Figure 8 and Figure 9 provide the complete ablation study results on $K$ and $\epsilon$.
Figure 8. Ablation study of $K$ on $\mathsf{CARIBOU}$ (colored boxes) and GAP (blue lines). For each pair of $\epsilon$ and dataset (Chain-S, Chain-M and Chain-L), different $K$ are used.
Figure 9. Ablation study of $\epsilon$ on $\mathsf{CARIBOU}$ (colored boxes) and GAP (blue lines). For each pair of $K$ and dataset (Chain-S, Chain-M and Chain-L), different $\epsilon$ are used.
E.4. Choices of Hyper-parameters
Appendix F More Related Works
Long-range graphs and GNNs.
Despite GNNs’ success in structured data, modeling long-range dependencies remains challenging due to over-smoothing (Chen et al., 2020a), vanishing gradients, and exponentially growing receptive fields. Enforcing DP in long-range GNNs is particularly challenging, as these models aggregate node embeddings over multiple hops and broader neighborhoods. Accordingly, injecting DP noise must account for multi-hop aggregation, requiring careful design to balance privacy and utility.
Differential privacy for graph structures.
Euclidean data clearly states which data points are associated with a particular individual, for example, tabular data. Unlike tabular data, graph association can be interpreted into the combination by edges and nodes. Corresponding to instance-level DP, the “instance” of graph data can be an edge or a node, naturally called edge DP and node DP. Early stage works (Gehrke et al., 2011; Mir and Wright, 2009; Wang et al., 2013; Chen et al., 2014; Lu and Miklau, 2014) started by private statistics estimation or counting, or release of private graphs under edge DP. Node DP protects against revealing the presence or absence of an individual node along with all its adjacent edges. Prior works consider various node-privacy algorithms (Kasiviswanathan et al., 2013; Iftikhar and Wang, 2021; Day et al., 2016; Xia et al., 2021; Imola et al., 2021) tailored to specific graph statistics constrained by bounded node degrees. Node DP typically offers a stronger level of protection compared to edge privacy, as inclusion of a node implies the inclusion of some particular edges. Node DP is also considered harder to achieve than edge DP if maintaining reasonable utility loss.
Differentially private GNNs.
Sharing model updates can lead to privacy risks, as adversaries may reconstruct training data. Such risks (Carlini et al., 2022; Ye et al., 2022) stem from overfitting, where neural networks memorize training data (Zhang et al., 2021). Graph structures, which encode relational information, are widely used in applications like intrusion detection (Apruzzese et al., 2023; Sommer and Paxson, 2010), social recommendation (Xia et al., 2023), and drug discovery (Zhong and Mottin, 2023). Graph neural networks (GNNs) have emerged as a key approach for learning over graph-structured data, but their message-passing mechanisms can leak sensitive information about nodes and their neighbors. To address these risks, several works (Sajadmanesh and Gatica-Perez, 2020, 2021; Sajadmanesh et al., 2023; Gupta et al., 2012) propose differentially private GNNs (Daigavane et al., 2021; Olatunji et al., 2021a; Sajadmanesh et al., 2023). This necessitates injecting calibrated noise after each message passing layer, where the noise scale is proportional to $K$ for a fixed privacy budget $\epsilon$. For instance, Wu et al. (Wu et al., 2022a) achieve edge-level privacy by adding Laplace noise to adjacency matrix entries under a small DP budget. Kolluri et al. (Kolluri et al., 2022) improve privacy-utility trade-offs by decoupling graph structure from the neural network architecture, querying the graph only when necessary. However, leveraging contractive hidden states in graph learning for enhanced privacy analysis remains an underexplored avenue.
Privacy analysis for iterations.
Prior works on differentially private graph learning analyze cumulative privacy costs by composing iteration-wise privacy guarantees using calibrated noise and privacy composition theorems. These theorems enable modular analysis of complex algorithms by controlling the total privacy budget. Feldman et al. (Feldman et al., 2018, 2020) address privacy analysis for gradient computations over a single training epoch under smooth and convex loss functions. However, extending these guarantees to multiple epochs often relies on composition bounds, such as privacy amplification by subsampling. Recent advancements (Chourasia et al., 2021; Altschuler and Talwar, 2022; Altschuler et al., 2024) improve privacy analysis for convex and strongly convex losses by eliminating dependence on infinite iterations, leveraging the Rényi DP framework.
In the typical differentially private setting, each access to the graph data must be protected. For multi-hop message passing GNNs (e.g., GAP (Sajadmanesh et al., 2023)), the privacy guarantee scales linearly with the number of hops $K$. As $K$ increases, the enlarged noise scale and cumulative effect of noise can significantly degrade model utility—precisely when deeper architectures are needed to capture long-range dependencies. This creates a fundamental tension: models require sufficient depth to learn meaningful long-range interactions, yet each additional layer demands more privacy noise, eventually overwhelming the signal. Consequently, standard privacy mechanisms could fail to effectively learn in graph structures requiring multi-hop propagation, as the privacy-utility tradeoff becomes increasingly unfavorable with model depth. Despite progress, foundational questions about privacy costs remain challenging, even for smooth convex losses over bounded domains (Altschuler and Talwar, 2022). Recent works (Altschuler and Talwar, 2022; Ye and Shokri, 2022) reveal that privacy leakage does not increase indefinitely with iterations. They demonstrate that after a brief burn-in period, additional iterations of SGD do not significantly impact privacy.
Privacy Audit and Attacks.
The goal of privacy auditing (Steinke et al., 2023) is to empirically estimate the privacy leakage of an algorithm. The main techniques used for privacy auditing is typically mounting a membership inference attack (MIA) (Carlini et al., 2022; Olatunji et al., 2021b) that predicts the existence or absence of an individual example in the training set. In contrast to the upper bound obtained from theoretical privacy analysis, privacy auditing via MIA achieves the experimental true positive rate (TPR) and false positive rate (FPR) and then presents a lower bound on the privacy budget $\epsilon$.